### v2

# Transitivity in wreath products with symmetric groups

Abstract.
It is known that the notion of a transitive subgroup of a permutation group $P$ extends naturally to the subsets of $P$. We study transitive subsets of the wreath product $G\wr S_{n}$, where $G$ is a finite abelian group. This includes the hyperoctahedral group for $G=C_{2}$. We give structural characterisations of transitive subsets using the character theory of $G\wr S_{n}$ and interpret such subsets as designs in the conjugacy class association scheme of $G\wr S_{n}$. In particular, we prove a generalisation of the Livingstone-Wagner theorem and give explicit constructions of transitive sets. Moreover, we establish connections to orthogonal polynomials, namely the Charlier polynomials, and use them to study codes and designs in $C_{r}\wr S_{n}$. Many of our results extend results about the symmetric group $S_{n}$.
2020 Mathematics Subject Classification:
05B99, 05E30, 20C99
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 491392403 – TRR 358
1. Introduction
The study of groups with a transitive action on an interesting set of objects is a classical and well-investigated topic (see for example [5]). Two kinds of transitivity that have attracted substantial research are $t$-transitivity and $t$-homogeneity. A subgroup of the symmetric group $S_{n}$ is called $t$-transitive if it acts transitively on the tuples of $t$ pairwise distinct elements of $\{1,\ldots,n\}$. A subgroup of the symmetric group $S_{n}$ is called $t$-homogeneous if it acts transitively on the subsets of $\{1,\ldots,n\}$ of size $t$. It is immediate that a $t$-transitive subgroup is also $t$-homogeneous and $(t-1)$-transitive. Livingstone and Wagner proved the following famous result.
Theorem 1.1 ([17, Thm. 2]).
Let $G$ be a subgroup of $S_{n}$ that is $t$-homogeneous for some $t$ with $21\leq t\leq n/$ 2. Then $G$ is also $(t-1)$-homogeneous.
This theorem was generalised by Martin and Sagan [19] in multiple ways. Their results include subsets instead of only subgroups and apply to a more general form of transitivity called $\lambda$-transitivity where $\lambda$ is a partition of $n$. Let us briefly explain these notions.
Let $G$ be a group acting on a set $\Omega$. A subset $Y\subseteq G$ is called transitive on $\Omega$ if there exists a constant $c>0$ such that for any $a,b\in\Omega$, there are exactly $c$ elements $g\in Y$ with $ga=b$. If $c=1$, then $Y$ is called sharply transitive. For every $a\in\Omega$, the integer $c$ is precisely the number of elements in $Y$ that stabilise $a$, so $c=|Y\cap\operatorname{Stab}(a)|$ is independent of $a$. If $Y$ is a subgroup, then this definition coincides with the definition of a transitive group action.
One reason to study transitive sets that are not subgroups is their size. It is well-known that for $t\geq 6$, the only $t$-transitive groups are $S_{n}$ (for $n\geq 6$) and $A_{n}$ (for $n\geq 8$). However, it is possible to construct $t$-transitive subsets of $S_{n}$ with $t\geq 6$ that are considerably smaller than $S_{n}$ or $A_{n}$, see [19, Section 6] for details.
The results by Martin and Sagan also deal with transitivity on ordered set partitions of $\Omega$. An (integer) partition $\lambda$ of a non-negative integer $n$ is a sequence $\lambda=(\lambda_{1},\lambda_{2},\ldots)$ of non-negative integers with $\lambda_{i}\geq\lambda_{i+1}$ for every $i\in{\mathbb{N}}$ and $\sum_{i\geq 1}\lambda_{i}=n$. For two partitions $\lambda=(\lambda_{1},\lambda_{2},\ldots)$ and $\mu=(\mu_{1},\mu_{2},\ldots)$ of $n$, we write $\lambda\unlhd\mu$ if
$\sum_{i=1}^{k}\lambda_{i}\leq\sum_{i=1}^{k}\mu_{i}\quad\text{ for all }k\geq 1.$
An ordered set partition $P=(P_{1},P_{2},\ldots)$ of $\{1,2,\ldots,n\}$ is a tuple of pairwise disjoint subsets $P_{i}\subseteq\{1,\ldots,n\}$ whose union is $\{1,\ldots,n\}$. For an integer partition $\lambda$, such an ordered set partition is called a $\lambda$-partition if $|P_{i}|=\lambda_{i}$ for all $i\geq 1$. A subset $Y$ of $S_{n}$ is called $\lambda$-transitive if it is transitive on the set of all $\lambda$-partitions. One of the main theorems in [19] is the following.
Theorem 1.2 ([19, Thm. 3]).
Let $\lambda$ and $\mu$ be partitions of $n$ and let $Y$ be a $\lambda$-transitive subset of $S_{n}$. If $\lambda\unlhd\mu$, then $Y$ is also $\mu$-transitive.
Since $t$-homogeneity is the same as $\lambda$-transitivity for the partition $\lambda=(n-t,t)$, Theorem 1.2 can be applied to $t$-homogeneity. Taking $t\in{\mathbb{N}}$ with $21\leq t\leq n/$ 2, $\lambda=(n-t,t)$ and $\mu=(n-t+1,t-1)$, we find that $\lambda\unlhd\mu$, so Theorem 1.1 is an immediate corollary of Theorem 1.2.
In this paper, we extend these results to wreath products with the symmetric group $S_{n}$. These products can be interpreted as groups of permutations of the numbers $1$ to $n$ where every number is coloured with one of $r$ different colours for some $r\in{\mathbb{N}}$. More precisely, they act on the set $\{(c,i):c\in\{1,\ldots,r\},i\in\{1,\ldots,n\}\}$. The action of the wreath product is imprimitive and for each $i\in[n]$, the sets $\{(c,i):c\in\{1,\ldots,r\}\}$ form blocks of imprimitivity. Thus, it is not possible for such a group to be 2-transitive, but there are many other interesting types of transitivity that we define and investigate in this paper. We show that these types of transitivity are meaningful in that they satisfy a version of the Livingstone-Wagner theorem and that they can be characterised algebraically. The class of wreath products with symmetric groups also includes symmetry groups of regular polytopes, which gives rise to a geometric interpretation of transitivity that shall motivate our studies and will serve as a basis for further generalisation.
Consider an $n$-dimensional regular polytope, for example a cube in the case $n=3$ (see [7] for an introduction to regular polytopes). A flag is a chain $F_{0}\subseteq F_{1}\subseteq\ldots\subseteq F_{n}$ of parts of the polytope, where each $F_{i}$ is an $i$-dimensional face of the polytope (so $\dim F_{i}-\dim F_{i-1}=1$ for every $i$). By definition, the symmetry group of a regular polytope acts transitively on its flags. It is easy to see that for a regular polytope, the symmetry group also acts transitively on the parts of the polytope, for example the set of edges. We are interested in algebraic characterisations of sets that are transitive on substructures of the polytope and how they can be compared to each other.
In order to illustrate this, let us consider the $n$-dimensional cube. Let $k$ be an integer with $0\leq k\leq n$ and $\sigma=(\sigma_{0},\sigma_{1},\ldots)$ be a sequence of positive integers that sum to $n-k$ (so $\sigma$ is a composition of $n-k$, not necessarily a partition of $n-k$). We call a chain of faces $F_{0}\subseteq F_{1}\subseteq\ldots$ of the cube a $(\sigma,k)$-flag if $\dim F_{0}=k$ and $\dim F_{i}-\dim F_{i-1}=\sigma_{i}$ for every $i$. For example, an $((n-1),1)$-flag of the cube is simply an edge. A $((2,n-2),0)$-flag is a pair of a vertex $V$ and a 2-dimensional face $F$ such that $V\subseteq F$, while an $((n-2,2),0)$-flag is a pair of a vertex $V$ and an $(n-2)$-dimensional face $F$ such that $V\subseteq F$. We call a subset $Y$ of the symmetry group of the cube $(\sigma,k)$-transitive if it is transitive on the set of all $(\sigma,k)$-flags. We shall later see that the order of the parts of $\sigma$ does not matter. That is, a subset $Y$ is $(\sigma,k)$-transitive if and only if it is $(\widetilde{\sigma},k)$-transitive where $\widetilde{\sigma}$ is a permutation of $\sigma$. Different permutations $\widetilde{\sigma}$ of $\sigma$ result in different geometric interpretations of $(\sigma,k)$-transitivity, but all of them are equivalent. Thus, for simplicity, we will assume that $\sigma$ is a partition of $n-k$.
The following is a special case of Theorem 5.7, one of our main results. In the theorem, $\sigma\cup(k)$ denotes the partition obtained by adding a part of size $k$ to $\sigma$. In the case $k=0$, we have $\sigma\cup(k)=\sigma$.
Theorem 1.3.
Let $n\in{\mathbb{N}}$ and let $k$ and $l$ be non-negative integers with $k,l\leq n$. Let $\sigma$ be a partition of $n-k$, let $\tau$ be a partition of $n-\ell$ and let $Y$ be a $(\sigma,k)$-transitive subset of the symmetry group of the $n$-dimensional cube. If $k\leq\ell$ and $(\sigma\cup(k))\unlhd(\tau\cup(\ell))$, then $Y$ is also $(\tau,\ell)$-transitive.
As an example, we can take $n=3$, $k=1$, $\ell=2$, $\sigma=(2)$ and $\tau=(1)$ in Theorem 1.3 to obtain the following statement: If a subset $Y$ of the symmetry group of the 3-dimensional cube is transitive on the set of its 12 edges, then it is also transitive on the set of its 6 faces.
We obtain our results by studying association schemes of the symmetry groups of regular polytopes. This leads us to studying Coxeter groups and wreath products. In what follows, it will be helpful to keep the basic example of the cube in mind.
We organise this paper in the following way: In Section 2, we describe the wreath products we are interested in. We also recall some basic facts about association schemes. Moreover, we describe a way to investigate transitive subsets of a group $G$ by studying an association scheme associated with it. This is done using representation theory. Section 3 then deals with the representation theory of the groups we want to investigate. In Section 4, we combine the results from Sections 2 and 3 to show that the transitive sets we are interested in can be characterised as special subsets in the corresponding association scheme. We then use this characterisation in Section 5 to describe relations between sets that are transitive on different sets of objects and obtain a generalisation of Theorem 1.2. In Section 6, we study Charlier polynomials and their connections to transitive sets in wreath products with symmetric groups, and in Section 7, we describe an explicit construction of transitive sets. Finally, we describe an application of our results to finite projective planes in Section 8.
2. Transitivity and association schemes
In this section, we fix some notations and introduce the main definitions of transitive sets and association schemes.
2.1. Transitive subsets
In this section, $(G,+)$ is a finite abelian group. The wreath product $W=G\wr S_{n}$ is a group, namely the semidirect product $W=G^{n}\rtimes S_{n}$ where $\sigma\in S_{n}$ acts on $G^{n}$ by mapping $(g_{1},\ldots,g_{n})\in G^{n}$ to $g^{\sigma}=(g_{\sigma^{-1}(1)},\ldots,g_{\sigma^{-1}(n)})$. We can represent every element of $W$ as a pair $(g,\sigma)$ with $g\in G^{n}$ and $\sigma\in S_{n}$. The product of two elements $(f,\pi),(g,\sigma)\in W$ is given by
$(f,\pi)(g,\sigma)=(f+g^{\pi},\pi\sigma).$
For more information about wreath products see [9, Chapter 2.6], for example.
If $G=C_{1}$ or $G=C_{2}$, then $W$ is the Coxeter group of type $A_{n-1}$ or $B_{n}$, respectively. If $G=C_{r}$ is the cyclic group of order $r$, then $W$ is the complex reflection group $G(r,1,n)$. The group $C_{r}\wr S_{n}$ is also called the generalised symmetric group. For more information about Coxeter groups and complex reflection groups, we refer the reader to [13].
Given an integer $n\in{\mathbb{N}}$, we write $[n]$ for the set $\{1,\ldots,n\}$. There is an action of $W$ on the set $M=G\times[n]$. For $((g_{1},\dots,g_{n}),\pi)\in W$ and $(c,i)\in G\times[n]$, it is given by
$((g_{1},\dots,g_{n}),\pi)\cdot(c,i)=(c+g_{\pi(i)},\pi(i)).$
It is easy to see that the wreath product $W$ consists of all permutations $\sigma$ of the set $M$ such that
$\sigma(0,i)=(c,j)\;\Longrightarrow\;\sigma(g,i)=(g+c,j)$
for all $g\in G$ and all $i\in[n]$ together with composition of mappings. We can think of the set $M$ as the set of $|G|$ copies of $[n]$ where each $i\in[n]$ appears in $|G|$ different colours.
In this paper, we study the action of $W$ on certain sets $\Omega$. We first collect some basic results about transitive group actions. They are easily found in the literature, see for example [5] for details.
Lemma 2.1.
Let $W$ be a group acting transitively on a set $\Omega$.
  1. (a)
    For every $x,y\in\Omega$ we have
    $\operatorname{Stab}(x)=w\operatorname{Stab}(y)w^{-1}$
    for some $w\in W$.
  2. (b)
    The action of $W$ on $\Omega$ is equivalent to the action of $W$ by left multiplication on the cosets $W/H=\{wH\mid w\in W\}$ where $H=\operatorname{Stab}(x)$ for some $x\in\Omega$.
  3. (c)
    Two coset actions on $W/H$ and $W/K$ by left multiplication are isomorphic if and only if $H$ and $K$ are conjugate.
This lemma allows us to investigate transitivity solely within the group $W$ itself. Every subgroup $H$ of $W$ gives rise to a transitive action of $W$ on the cosets $W/H$ and every transitive action arises this way. We can think of each subgroup $H$ as a certain type of transitivity.
We aim to investigate as many different transitivity types in the group $W$ as possible. In order to illustrate this point of view, let us look at $t$-transitivity. The action of $W$ on $M$ is imprimitive, so no subset of $W$ under this action is ever $t$-transitive for $t\geq 2$. However, we can consider a different definition of $t$-transitivity, namely transitivity on $t$-tuples where each element is from a different block. Let
$\Omega=\left\{\big{(}(g_{1},a_{1}),\ldots,(g_{t},a_{t})\big{)}\in M^{t}:g_{i}\in G,\,a_{i}\in[n]\text{ with }a_{i}\neq a_{j}\text{ for }i\neq j\right\}$
be the set of all $t$-tuples of elements of $M$ where the elements of $[n]$ are pairwise distinct. Then the stabiliser of an element $x\in\Omega$ is the subgroup
$\text{Stab}(x)=G\wr S_{A}\;\cong\;G\wr S_{n-t}=G^{n-t}\rtimes S_{n-t},$
where $A=[n]\,\setminus\,\{a_{1},\ldots,a_{t}\}$. Thus, in view of Lemma 2.1, a subset $Y$ of $W$ is transitive on $\Omega$ if and only if it is transitive on the cosets $W/H$ where $H=G\wr S_{n-t}$.
2.2. Characters
Since there is a close connection between association schemes of groups and their representations, we now turn to representation and character theory. We refer the reader to [22] for a more extensive introduction to the representation theory of finite groups.
Let $G$ be a finite group. A (complex) representation of $G$ is a homomorphism $M:G\to\operatorname{GL}_{n}({\mathbb{C}})$ where $\operatorname{GL}_{n}({\mathbb{C}})$ denotes the group of invertible $(n\times n)$-matrices over ${\mathbb{C}}$. Each representation $M$ has a character $\chi:G\to{\mathbb{C}}$ which is given by $\chi(g)=\operatorname{tr}M(g)$ where $\operatorname{tr}$ denotes the trace function. The degree $\deg(\chi)$ of a character $\chi$ is the dimension of the vector space of the underlying representation. It is also given by $\deg(\chi)=\chi(e)=\operatorname{tr}I_{n}=n$ where $e$ is the identity of $G$ and $I_{n}$ is the identity matrix. Every representation can be decomposed into so called irreducible representations, and hence every character can be expressed as a sum of irreducible characters.
Functions from $G$ to ${\mathbb{C}}$ that are constant on conjugacy classes are called class functions. Since all characters are class functions, we can write $\chi(C)$ for the value of the character $\chi$ on the conjugacy class $C$ of $G$. The class functions form a vector space over ${\mathbb{C}}$ that is generated by the irreducible characters. It has an inner product given by
$\langle\chi,\psi\rangle=\frac{1}{|G|}\sum_{g\in G}\chi(g)\overline{\psi(g)}$
[…]
If $G$ acts on a set $\Omega$, then the function $\xi_{\Omega}:G\to{\mathbb{C}}$ with $\xi_{\Omega}(g)=|\operatorname{Fix}(g)|$ is called the permutation character of $G$ on $\Omega$. If $H\subseteq G$ is a subgroup, then $G$ acts transitively on the set of left cosets $G/H$ via left multiplication and the corresponding permutation character $\xi_{H}$ is called the permutation character of $H$. By Lemma 2.1, the permutation character of a transitive action of $G$ on $\Omega$ equals the permutation character $\xi_{H}$ where $H$ is the stabiliser of some element $\omega\in\Omega$.
Let $H\subseteq G$ be a subgroup. If $\chi$ is a character of $G$, then restricting $\chi$ to $H$ gives a character of $H$. This character is called the restriction of $\chi$ to $H$ and denoted by $\chi\downarrow^{G}_{H}$. There is a way to extend a character of a subgroup $H$ to a character of $G$ called induction. If $\chi$ is a character of $H$, then the induced character is denoted by $\chi\uparrow^{G}_{H}$.
The function $1_{G}$ with $11_{G}(g)=$ 1 for all $g\in G$ is a character of $G$, called the trivial character or the principal character of $G$. If we restrict $1_{G}$ to a subgroup $H\subseteq G$, then we get the trivial character of $H$, so $1_{G}\downarrow^{G}_{H}=1_{H}$. It is well-known that the induction of the trivial character $1_{H}$ of $H$ to $G$ equals the permutation character of $H$, that is
$1_{H}\uparrow^{G}_{H}=\xi_{H}.$
In order to prove our results about permutation characters, we need Mackey’s formula. It describes the result of inducing a character followed by restriction. Let $H,K\subseteq G$ be subgroups. If $\chi$ is a character of $H$ and $s\in G$, we define a character $\chi^{s}$ of $H^{s}=sHs^{-1}$ via
$\chi^{s}(x)=\chi(s^{-1}xs)$
for every $x\in H^{s}$. We have the following theorem (see for example [22, Section 7.3]).
Theorem 2.2 (Mackey).
With the notation as above, we have
$\chi\uparrow^{G}_{H}\downarrow^{G}_{K}=\sum_{s\in R}\chi^{s}\downarrow^{H^{s}}_{H^{s}\cap K}\uparrow^{K}_{H^{s}\cap K}$
where $R$ is a system of representatives for the double cosets $KgH$ with $g\in G$.
2.3. The conjugacy class association scheme
We will mainly study the combinatorics of $G\wr S_{n}$ from the viewpoint of association schemes. We refer the reader to [1] for an extensive introduction to association schemes. Every finite group gives rise to an association scheme (see [1, Section 2.7] for details), called the conjugacy class scheme of the group, but the theory of association schemes is much more general than that. We collect the relevant details about conjugacy class schemes here. In view of Section 2.4, we will first describe conjugacy class schemes in general and later specialise to the group $G\wr S_{n}$.
For finite and non-empty sets $X$ and $Y$, let ${\mathbb{C}}(X,Y)$ denote the set of all complex $|X|\times|Y|$-matrices with rows indexed by $X$ and columns indexed by $Y$. For a matrix $A\in{\mathbb{C}}(X,Y)$ and $x\in X$, $y\in Y$, the $(x,y)$-entry of $A$ is denoted by $A(x,y)$. If $|Y|=1$, then we omit $Y$, so ${\mathbb{C}}(X)$ is the set of complex column vectors indexed by $X$.
Let $G$ be a finite group and let $C_{0}=\{e\},C_{1},\ldots,C_{n}$ denote its conjugacy classes. For $i\in\{0,\ldots,n\}$, let $D_{i}\in{\mathbb{C}}(G,G)$ be given by
(1) $D_{i}(x,y)=\begin{cases}1,&\text{ if }x^{-1}y\in C_{i}\\ 0,&\text{ otherwise}.\end{cases}$
Let $\mathbb{A}=\operatorname{sp}(D_{0},\ldots,D_{n})$ be the vector space generated by $D_{0},\ldots,D_{n}$ over the complex numbers. Then $\mathbb{A}$ is a commutative matrix algebra that contains the identity matrix and is closed under conjugate transposition. Thus, the zero-one-matrices $D_{i}$ define an association scheme, called the conjugacy class scheme of $G$. The algebra $\mathbb{A}$ is called the Bose-Mesner-algebra of this association scheme.
Since $\mathbb{A}$ is commutative, all of its elements can be simultaneously diagonalised, and therefore there exists a basis $E_{0},\ldots,E_{n}$ of $\mathbb{A}$ consisting of Hermitian matrices with the property
(2) $E_{k}E_{l}=\delta_{kl}E_{k}.$
The matrices $E_{k}$ are the minimal idempotents of $\mathbb{A}$. These matrices are given by (see [1, Theorem II.7.2])
$E_{k}=\frac{\deg(\chi_{k})}{|G|}\sum_{l=0}^{n}\chi_{k}(C_{l})D_{l}.$
Using (1), the entries of $E_{k}$ are given by
(3) $E_{k}(x,y)=\frac{\deg(\chi_{k})}{|G|}\chi_{k}(x^{-1}y).$
Let $V_{k}$ be the column space of $E_{k}$. The spaces $V_{k}$ are called the eigenspaces of the association scheme because they are the common eigenspaces of the matrices $D_{i}$. From (2) we obtain that these vector spaces are pairwise orthogonal and that
(4) ${\mathbb{C}}(G)=\bigoplus_{k=0}^{n}V_{k}.$
Let $Y$ be a non-empty subset of $G$. We can associate two sequences of numbers to $Y$: The inner distribution $a$ and the dual distribution $a^{\prime}$. The inner distribution of $Y$ is the tuple $(a_{0},\ldots,a_{n})$, where
(5) $a_{i}=\frac{1}{|Y|}\sum_{x,y\in Y}D_{i}(x,y),$
and the dual distribution of $Y$ is the tuple $(a^{\prime}_{0},\ldots,a^{\prime}_{n})$, where
(6) $a^{\prime}_{k}=\frac{|G|}{|Y|}\sum_{x,y\in Y}E_{k}(x,y).$
We can use (3) to obtain a formula for the dual distribution $a^{\prime}$ in terms of characters:
(7) $a^{\prime}_{k}=\frac{\deg(\chi_{k})}{|Y|}\sum_{x,y\in Y}\chi_{k}(x^{-1}y)$
It is immediate that the inner distribution is non-negative. The same holds for the dual distribution. Let $\mathbbm{1}_{Y}\in{\mathbb{C}}(G)$ be the characteristic vector of $Y$, so $\mathbbm{1}_{Y}(x)=1$ if $x\in Y$ and $\mathbbm{1}_{Y}(x)=0$ otherwise. Since the matrix $E_{k}$ is hermitian, it follows from (7) that
(8) $\frac{|Y|}{|G|}a^{\prime}_{k}=\mathbbm{1}_{Y}^{\top}E_{k}\mathbbm{1}_{Y}=\mathbbm{1}_{Y}^{*}E_{k}^{*}E_{k}\mathbbm{1}_{Y}=\lVert E_{k}\mathbbm{1}_{Y}\rVert^{2}.$
Thus, the dual distribution is real and non-negative. Furthermore, the case $a^{\prime}_{k}=0$ occurs if and only if $\mathbbm{1}_{Y}$ is orthogonal to $V_{k}$.
The main reason why the theory of association schemes is such a powerful tool in combinatorics, is the observation that interesting combinatorial structures can often be characterised using the inner or dual distribution. Delsarte [8] calls these objects cliques and designs, respectively.
Definition 2.3.
Let $Y$ be a non-empty subset of $G$ with inner distribution $(a_{0},\ldots,a_{n})$ and dual distribution $(a^{\prime}_{0},\ldots,a^{\prime}_{n})$.
  1. (a)
    Let $D\subseteq[n]$. We call $Y$ a $D$-clique or a $D$-code if $a_{d}=0$ for every $d\not\in D$.
  2. (b)
    Let $T\subseteq[n]$. We call $Y$ a $T$-design if $a^{\prime}_{t}=0$ for every $t\in T$.
We show in Section 4 that the structures we are interested in are in fact $T$-designs in the sense of Delsarte.
2.4. The general method
We now describe the method we use to connect transitive subsets to designs in an association scheme. The technique can be applied to any finite group $G$ acting on a set $\Omega$. The result is essentially that if the decomposition of the permutation character on $\Omega$ is known, then this decomposition gives rise to a characterisation of transitivity on $\Omega$ as a $T$-design in an association scheme.
Let $G$ be a finite group acting on a set $\Omega$. We define the incidence matrix of this action to be the matrix $M\in{\mathbb{C}}(G,\Omega\times\Omega)$ given by
$M(g,(a,b))=\begin{cases}1&\text{if $ga=b$},\\ 0&\text{otherwise}.\end{cases}$
Denote by $\operatorname{colsp}(M)$ the vector space generated by the columns of $M$. The following theorem is crucial.
Theorem 2.4.
Let $G$ be a finite group acting on a set $\Omega$ with incidence matrix $M$. Let $\chi_{0},\chi_{1},\dots,\chi_{n}$ be the (complex) irreducible characters of $G$ and let $V_{0},V_{1},\dots,V_{n}$ be the corresponding eigenspaces of the conjugacy class association scheme of $G$. Define $I\subseteq\{0,1,\dots,n\}$ such that $k\in I$ if and only if $\chi_{k}$ is an irreducible constituent of the permutation character $\xi$ of $G$ on $\Omega$. Then
$\operatorname{colsp}(M)=\bigoplus_{k\in I}V_{k}.$
Proof.
We have $MM^{\top}\in{\mathbb{C}}(G,G)$ and for $g,h\in G$ we have
$(MM^{\top})(g,h)=\sum_{(a,b)\in\Omega\times\Omega}M(g,(a,b))M(h,(a,b)).$
The summands are either 0 or 1. A summand is equal to 1 if and only if $ga=b$ and $ha=b$, so $(g^{-1}h)a=a$. Thus, the entry of $MM^{\top}$ at position $(g,h)$ is equal to the number of fixed points $|\operatorname{Fix}(g^{-1}h)|$ of $g^{-1}h$.
For the permutation character $\xi$, we have $\xi(g)=|\operatorname{Fix}(g)|$. Thus, considering the matrix $P\in{\mathbb{C}}(G,G)$ corresponding to $\xi$ given by $P(g,h)=\xi(g^{-1}h)=|\operatorname{Fix}(g^{-1}h)|$, we obtain $P=MM^{\top}$. This implies $\operatorname{colsp}(M)=\operatorname{colsp}(MM^{*})=\operatorname{colsp}(MM^{\top})=\operatorname{colsp}(P)$.
Now consider the decomposition
$\xi=m_{0}\chi_{0}+\ldots+m_{n}\chi_{n}$
of $\xi$ into irreducible characters $\chi_{k}$ where $m_{k}\in{\mathbb{N}}_{0}$. Recall that by (3), the idempotent matrix $E_{k}$ of the conjugacy class association scheme is given by
$E_{k}(g,h)=c_{k}\chi_{k}(g^{-1}h)$
for some constant $c_{k}>0$. Since
$P(g,h)=\xi(g^{-1}h)=\sum_{k=0}^{n}m_{k}\chi_{k}(g^{-1}h)=\left(\sum_{k=0}^{n}m_{k}c^{-1}_{k}E_{k}\right)(g,h),$
we obtain (since $k\in I$ if and only if $m_{k}\neq 0$)
(9) $P=\sum_{k=0}^{n}m_{k}c^{-1}_{k}E_{k}=\sum_{k\in I}m_{k}c^{-1}_{k}E_{k}.$
Recall that $\operatorname{colsp}(E_{k})=V_{k}$ and that the spaces $V_{k}$ are pairwise orthogonal. Hence, the column space of $P$ is contained in $\bigoplus_{k\in I}V_{k}$. Conversely, let $k\in I$ and $v$ be a column vector of $E_{k}$. From (2), we have $E_{k}v=v$ and $E_{l}v=0$ for $l\neq k$, so it follows from (9) that
$Pv=PE_{k}v=m_{k}c_{k}^{-1}v.$
Since $m_{k}\neq 0$ for $k\in I$, we conclude that $V_{k}=\operatorname{colsp}(E_{k})\subseteq\operatorname{colsp}(P)$ for every $k\in I$. Thus, we have
$\operatorname{colsp}(M)=\operatorname{colsp}(P)=\bigoplus_{k\in I}V_{k}.\qed$
Now we can describe the connection between transitivity and the dual distribution.
Theorem 2.5.
With the same notation as in Theorem 2.4, let $Y$ be a non-empty subset of $G$ with dual distribution $(a^{\prime}_{0},a^{\prime}_{1},\dots,a^{\prime}_{n})$. If $G$ acts transitively on $\Omega$, then $Y$ is transitive on $\Omega$ if and only if
$a^{\prime}_{k}=0\quad\text{for each $k\in I\,\setminus\{0\}$}$
where $I$ is the set of all $k\in\{0,\ldots,n\}$ such that $\chi_{k}$ is an irreducible constituent of the permutation character $\xi$ on $\Omega$.
Proof.
Note that $G$ acts transitively on $\Omega$ if and only if
$M^{\top}\mathbbm{1}_{G}=c\cdot\mathbbm{1}_{\Omega\times\Omega}$
for some constant $c$, since this equation means that for any pair $(\omega,\omega^{\prime})\in\Omega\times\Omega$ there exist precisely $c$ elements of $G$ that map $\omega$ to $\omega^{\prime}$. Thus, $Y$ is transitive on $\Omega$ if and only if
$M^{\top}\mathbbm{1}_{Y}=c^{\prime}\cdot\mathbbm{1}_{\Omega\times\Omega}$
for some constant $c^{\prime}$. This constant can be calculated by observing that if $Y$ is transitive, then so are the cosets $gY$ for $g\in G$. Since the disjoint union of transitive sets is a transitive set, we find that $c=|G/Y|\cdot c^{\prime}$ which gives
$c^{\prime}=\frac{|Y|}{|G|}\cdot c.$
Now we see that $Y$ is transitive on $\Omega$ if and only if
$\frac{1}{\lvert Y\rvert}M^{\top}\mathbbm{1}_{Y}=\frac{1}{\lvert G\rvert}M^{\top}\mathbbm{1}_{G},$
hence if and only if
$\mathbbm{1}_{Y}-\frac{\lvert Y\rvert}{\lvert G\rvert}\mathbbm{1}_{G}$
is orthogonal to the column space of $M$. By Theorem 2.4, the column space of $M$ is given by $\operatorname{colsp}(M)=\bigoplus_{k\in I}V_{k}$. Hence, $\mathbbm{1}_{Y}-\frac{\lvert Y\rvert}{\lvert G\rvert}\mathbbm{1}_{G}$ is orthogonal to the column space of $M$ if and only if it is orthogonal to the spaces $V_{k}$ with $k\in I$.
First, note that $V_{0}$ is spanned by $\mathbbm{1}_{G}$ and that
$\langle\mathbbm{1}_{Y}-\frac{\lvert Y\rvert}{\lvert G\rvert}\mathbbm{1}_{G},\mathbbm{1}_{G}\rangle=0.$
Now let $k\in\{1,\ldots,n\}$ and $v\in V_{k}$. Since the spaces $V_{k}$ are pairwise orthogonal and $\mathbbm{1}_{G}\in V_{0}$, we find that
$\langle\mathbbm{1}_{Y}-\frac{\lvert Y\rvert}{\lvert G\rvert}\mathbbm{1}_{G},v\rangle=\langle\mathbbm{1}_{Y},v\rangle.$
Thus, for $k\neq 0$, $\mathbbm{1}_{Y}-\frac{\lvert Y\rvert}{\lvert G\rvert}\mathbbm{1}_{G}$ is orthogonal to $V_{k}$ if and only if $\mathbbm{1}_{Y}$ is orthogonal to $V_{k}$. Recall that by (8), $\mathbbm{1}_{Y}$ is orthogonal to $V_{k}$ if and only if $a^{\prime}_{k}=0$, which concludes the proof. ∎
In view of Theorem 2.5, our next task is to study the representation theory of $G\wr S_{n}$ and decompose the permutation characters $\xi_{H}$ for a given subgroup $H$.
3. Wreath products with symmetric groups
In this section, we collect the definitions and results about permutation groups and their representation theory that we need later on. We mostly follow [18, Chapter I].
3.1. Partitions
An (integer) partition is a sequence $\lambda=(\lambda_{1},\lambda_{2},\ldots)$ of non-negative integers that sum to a finite number and satisfy $\lambda_{i}\geq\lambda_{i+1}$ for every $i\in{\mathbb{N}}$. The size of $\lambda$ is the number $|\lambda|=\sum_{i\geq 1}\lambda_{i}$ and if $|\lambda|=n$, we say that $\lambda$ is a partition of $n$. We often write partitions as finite sequences $(\lambda_{1},\ldots,\lambda_{k})$ only writing down the non-zero parts. We also write $k^{t}$ to denote $t$ parts of size $k$, for example $2^{1}1^{2}$ is the sequence $(2,1,1,0,0,\ldots)$. Let $\operatorname{Par}$ denote the set of all integer partitions (of any size). The unique partition of 00 will be denoted by $\emptyset$.
If $\lambda$ is a partition of $n$, then the Young diagram of $\lambda$ is an array of $n$ boxes with left-justified rows and top-justified columns such that row $i$ contains exactly $\lambda_{i}$ boxes. Each partition gives rise to a conjugate partition $\lambda^{\prime}$ whose parts are the number of boxes in the columns of the Young diagram of $\lambda$ (so the Young diagram of $\lambda^{\prime}$ is the transpose of the Young diagram of $\lambda$).
[…]
$\sum_{i=1}^{k}\lambda_{i}\leq\sum_{i=1}^{k}\mu_{i}\quad\text{ for all }k\geq 1.$
Note that if $|\lambda|=|\mu|$, then we have
$\lambda\unlhd\mu\;\Longleftrightarrow\;\lambda^{\prime}\unrhd\mu^{\prime}.$
Another helpful fact is that we have $\lambda\unlhd\mu$ if and only if the Young diagram of $\lambda$ can be transformed into the Young diagram of $\mu$ by moving boxes from the end of a row of $\lambda$ to a higher row one at a time. When interpreting a partition as a sequence, this corresponds to adding vectors of the form $(0,\ldots,0,1,0,\ldots,0,-1,0,\ldots)$ to the partition.
We use two methods for combining partitions. If $\lambda$ and $\mu$ are partitions, then $\lambda\cup\mu$ is the partition of $|\lambda|+|\mu|$ that has as its parts exactly the parts of $\lambda$ and $\mu$, and $\lambda+\mu$ is the componentwise sum, that is $\lambda+\mu=(\lambda_{1}+\mu_{1},\lambda_{2}+\mu_{2},\ldots)$. It holds that
$\left(\lambda\cup\mu\right)^{\prime}=\lambda^{\prime}+\mu^{\prime}\quad\text{ and }\quad\left(\lambda+\mu\right)^{\prime}=\lambda^{\prime}\cup\mu^{\prime}.$
3.2. Conjugacy classes
We now describe the conjugacy classes of $G\wr S_{n}=G^{n}\rtimes S_{n}$, where $G$ is finite abelian. We refer to [18, Appendix B, Chapter I] for further reading. The conjugacy classes of $G\wr S_{n}$ are indexed by functions ${\underline{\lambda}}:G\to\operatorname{Par}$ such that
$\sum_{g\in G}|{\underline{\lambda}}(g)|=n.$
We write
$\Lambda_{n}(G)=\{{\underline{\lambda}}:G\to\operatorname{Par}:\sum_{g\in G}|{\underline{\lambda}}(g)|=n\}$
for the set of all these partition valued functions ${\underline{\lambda}}$.
Consider an element $((g_{1},\dots,g_{n}),\pi)\in G\wr S_{n}$. The permutation $\pi\in S_{n}$ can be written as a product of disjoint cycles of the form $(i_{1},\ldots,i_{k})$. It turns out that the element $g_{i_{1}}+\ldots+g_{i_{k}}$ is determined up to conjugacy. Thus, since $G$ is abelian in our setting, it is determined uniquely. We call this element $g_{i_{1}}+\ldots+g_{i_{k}}$ the sign of the cycle $(i_{1},\ldots,i_{k})$. For each $g\in G$, we can determine all the cycles in $\pi$ that have sign $g$. Writing down all the cycle lengths that occur and ordering them decreasingly, we obtain a partition associated with the element $g$. The cycle type of an element $((g_{1},\dots,g_{n}),\pi)\in G\wr S_{n}$ is defined to be the mapping ${\underline{\lambda}}:G\to\operatorname{Par}$, where ${\underline{\lambda}}(g)$ equals the partition of cycle lengths of $\pi$ that have sign $g$. Then two elements of $G\wr S_{n}$ are conjugate if and only if they have the same cycle type. Since the sum of all cycle lengths of $\pi$ is $n$, the conjugacy classes of $G\wr S_{n}$ are indexed by the set $\Lambda_{n}(G)$.
We denote by $C_{\underline{\lambda}}$ the conjugacy class indexed by ${\underline{\lambda}}$. For example, the conjugacy class of the identity of $G\wr S_{n}$ is indexed by the element ${\underline{\lambda}}\in\Lambda_{n}(G)$ with ${\underline{\lambda}}(0)=(1^{n})$ and ${\underline{\lambda}}(g)=\emptyset$ for every $g\neq 0$.
3.3. Characters of wreath products
In this section, we review the representation theory of wreath products of the form $G\wr S_{n}=G^{n}\rtimes S_{n}$, where $G$ is a finite abelian group. We first consider the more general case $A\rtimes H$, where $A$ and $H$ are finite groups with $A$ being abelian. We essentially follow [22, Sect. 8.2].
Since $A$ is abelian, each irreducible character of $A$ has degree $1$ and the irreducible characters $\theta$ of $A$ form a group $A^{\prime}$. Note that $A$ is normal in $A\rtimes H$, so the group $H$ acts on $A^{\prime}$ by
$(h\theta)(a)=\theta(h^{-1}ah)\quad\text{for $h\in H$, $\theta\in A^{\prime}$ and $a\in A$}.$
Let $\theta$ be an irreducible character of $A$ and let $K$ be the stabiliser of $\theta$ under the action of $H$. Now let $\psi$ be an irreducible character of $K$. Then we can define the character $\theta\cdot\psi$ of $A\rtimes K$ by
$(\theta\cdot\psi)(ak)=\theta(a)\psi(k)\quad\text{for all $a\in A$, $k\in K$}.$
Finally, we induce $\theta\cdot\psi$ to a character $\chi_{\theta,\psi}$ of $A\rtimes H$, so that
$\chi_{\theta,\psi}=(\theta\cdot\psi)\uparrow_{A\rtimes K}^{A\rtimes H}.$
We have the following theorem.
Theorem 3.1 (Serre [22, Prop. 25]).
With the notation as above, $\chi_{\theta,\psi}$ is an irreducible character of $A\rtimes H$ and every irreducible character of $A\rtimes H$ is of this form. Moreover, if $\chi_{\theta,\psi}=\chi_{\theta^{\prime},\psi^{\prime}}$, then $\psi=\psi^{\prime}$ and $\theta$ and $\theta^{\prime}$ are in the same orbit under the action of $H$ on $A^{\prime}$.
We now want to apply Theorem 3.1 to the group $G\wr S_{n}=G^{n}\rtimes S_{n}$, so we need the characters of $G^{n}$ and the characters of some subgroups of the symmetric group. Since $G$ is abelian, all representations of $G$ are one-dimensional and the group of the corresponding characters of $G$ is isomorphic to $G$ itself. We fix an isomorphism $g\mapsto\theta^{g}$ from $G$ to its character group. Then the irreducible characters of $G$ are the characters $\theta^{g}$ for $g\in G$.
The characters of $G^{n}$ are products of characters of $G$. Thus, they are indexed by the tuples $(g_{1},\ldots,g_{n})\in G^{n}$. The irreducible character $\theta^{(g_{1},\ldots,g_{n})}$ is given by
$\theta^{(g_{1},\ldots,g_{n})}(x_{1},\ldots,x_{n})=\prod_{i=1}^{n}\theta^{g_{i}}(x_{i})\;\text{ for every }(x_{1},\ldots,x_{n})\in G^{n}.$
Now consider an element $(g_{1},\ldots,g_{n})\in G^{n}$. The symmetric group $S_{n}$ acts on the characters $\theta^{(g_{1},\ldots,g_{n})}$ by permuting the entries of $(g_{1},\dots,g_{n})$. For $g\in G$, let $k(g)$ be the number of times $g$ occurs in $(g_{1},\dots,g_{n})$. Then $\sum_{g\in G}k(g)=n$ and the stabiliser of the character $\theta^{(g_{1},\ldots,g_{n})}$ under the action of $S_{n}$ is
$\operatorname{Stab}(\theta^{(g_{1},\ldots,g_{n})})=\prod_{g\in G}S_{k(g)}$
where $S_{k(g)}$ permutes the indices $i\in[n]$ with $g_{i}=g$. This is a Young subgroup of $S_{n}$. The irreducible characters of a factor $S_{k(g)}$ are indexed by the partitions of $k(g)$ and denoted by $\psi^{\mu}$ where $\mu$ is a partition. Now, the irreducible characters of $\operatorname{Stab}(\theta^{(g_{1},\ldots,g_{n})})$ are given by
$\psi^{\underline{\mu}}=\bigotimes_{g\in G}\psi^{{\underline{\mu}}(g)},$
where ${\underline{\mu}}:G\to\operatorname{Par}$ is a function such that ${\underline{\mu}}(g)$ is a partition of $k(g)$. Combining $\psi^{\underline{\mu}}$ with the character $\theta^{(g_{1},\ldots,g_{n})}$ as in Theorem 3.1 gives an irreducible character of $G\wr S_{n}$. Since we only need the orbit of $\theta^{(g_{1},\ldots,g_{n})}$, so the numbers $k(g)$ for $g\in G$, all information needed is encoded in ${\underline{\mu}}$ (because $k(g)=|{\underline{\mu}}(g)|$). Thus, the irreducible characters of $G\wr S_{n}$ are indexed by the set of mappings ${\underline{\lambda}}:G\to\operatorname{Par}$ such that
$\sum_{g\in G}\lvert{\underline{\lambda}}(g)\rvert=n.$
This is the set $\Lambda_{n}(G)$ that we have seen earlier. For ${\underline{\lambda}}\in\Lambda_{n}(G)$, the corresponding irreducible character is denoted by $\chi^{{\underline{\lambda}}}$. For example, the trivial character of $G\wr S_{n}$ is $\chi^{{\underline{\lambda}}}$, where ${\underline{\lambda}}\in\Lambda_{n}(G)$ is the unique element with ${\underline{\lambda}}(0)=(n)$ and ${\underline{\lambda}}(g)=\emptyset$ for every $g\neq 0$.
Next, we present a different way to write down the irreducible characters of $G\wr S_{n}$. For a character $\theta$ of $G$ and a character $\psi$ of $S_{k}$, we define $\theta\wr\psi$ to be the character of $G\wr S_{k}$ given by
$(\theta\wr\psi)((g_{1},\dots,g_{k}),\pi)=\theta(g_{1})\cdots\theta(g_{k})\,\psi(\pi).$
Now consider ${\underline{\lambda}}\in\Lambda_{n}(G)$. Then we have partitions ${\underline{\lambda}}(g)$ for $g\in G$. The character $\chi^{{\underline{\lambda}}}$ of $G\wr S_{n}$ is given by
(10) $\chi^{{\underline{\lambda}}}=\bigotimes_{g\in G}\left(\theta^{g}\wr\psi^{{\underline{\lambda}}(g)}\right)\uparrow_{G\wr S}^{G\wr S_{n}},$
where
(11) $S=\prod_{g\in G}S_{|{\underline{\lambda}}(g)|}$
is a Young subgroup of $S_{n}$ and $G\wr S$ denotes the subgroup $G^{n}\rtimes S$ of $G\wr S_{n}=G^{n}\rtimes S_{n}$. Note that we did not specify a tuple $(g_{1},\ldots,g_{n})\in G^{n}$ as before. The element ${\underline{\lambda}}\in\Lambda_{n}(G)$ defines a tuple $(g_{1},\ldots,g_{n})$ up to a permutation of the entries (the element $g\in G$ appears $|{\underline{\lambda}}(g)|$ times). Different choices of tuples lead to conjugate Young subgroups $S$ and ultimately to the same character $\chi^{{\underline{\lambda}}}$. In most cases, it is not important how exactly the group $S$ lies in the group $S_{n}$. For example, $S_{\{1,2\}}\times S_{\{3,4\}}$ and $S_{\{1,4\}}\times S_{\{2,3\}}$ are different subgroups of $S_{4}$, but they are conjugate to each other. The first subgroup corresponds to the ordered set partition ($(\{1,2\},\{3,4\})$ while the latter corresponds to the ordered set partition ($(\{1,4\},\{2,3\})$. Everytime we write down a Young subgroup $S$, it implicitly comes with an ordered set partition $P$. In our case, the parts of $P$ are indexed by the group $G$. The factor $S_{|{\underline{\lambda}}(g)|}$ permutes the set of indices $i\in[n]$ of the tuple $(g_{1},\ldots,g_{n})$ such that $g_{i}=g$. This set of indices is also the part of $P$ that is indexed by $g\in G$.
If ${\underline{\lambda}}(g)=\emptyset$ for some $g\in G$, the corresponding factor in the product in (10) or (11) is trivial. The reader should interpret equation (10) as first choosing for each $g\in G$ how often each character $\theta^{g}$ appears, then choosing a partition ${\underline{\lambda}}(g)$ of the corresponding size to get a character $\psi^{{\underline{\lambda}}(g)}$, and then combining all these characters. In view of Theorem 3.1, the $\chi^{\underline{\lambda}}$ are all the irreducible characters of $G\wr S_{n}$.
4. Notions of Transitivity
We continue to consider the group $W=G\wr S_{n}$ for a fixed abelian group $G$. We will now focus on the different types of transitivity in $W$, which means the different $\Omega$ in Theorem 2.5.
4.1. Transitivity types
Let $S(G)$ be the set of subgroups of $G$. For a mapping ${\underline{\sigma}}:S(G)\to\operatorname{Par}$, we define
$\lVert{\underline{\sigma}}\rVert=\sum_{U\leq G}\lvert{\underline{\sigma}}(U)\rvert.$
We write
$\Sigma_{n}(G)=\{{\underline{\sigma}}:S(G)\to\operatorname{Par}:\lVert{\underline{\sigma}}\rVert=n\}$
for the set of all mappings ${\underline{\sigma}}:S(G)\to\operatorname{Par}$ satisfying $\lVert{\underline{\sigma}}\rVert=n$.
Each ${\underline{\sigma}}\in\Sigma_{n}(G)$ can be represented by an ordered set of Young diagrams, one for each $U\in G$, such that the total number of boxes equals $n$. We define a ${\underline{\sigma}}$-tableau to be a filling of the boxes with pairs $(c,i)$, where $c$ is a coset of a subgroup $U$ of $G$ and $i\in[n]$ with the following property. All numbers $i\in[n]$ appearing in the pairs $(c,i)$ are pairwise distinct and if $(c,i)$ appears in the Young diagram of the subgroup $U$, then $c\in G/U$. Two ${\underline{\sigma}}$-tableaux are row-equivalent if the corresponding rows of the two tableaux contain the same elements. This induces an equivalence relation on ${\underline{\sigma}}$-tableaux. A ${\underline{\sigma}}$-tabloid is an equivalence class of this relation.
Note that $W$ acts on the set of ${\underline{\sigma}}$-tabloids in the natural way. We define a notion of transitivity using ${\underline{\sigma}}$-tabloids.
Definition 4.1.
A subset $Y$ of $W$ is called ${\underline{\sigma}}$-transitive if $Y$ is transitive on the set of ${\underline{\sigma}}$-tabloids.
The stabiliser of a ${\underline{\sigma}}$-tabloid is
$H_{{\underline{\sigma}}}=\prod_{U\leq G}\left(U\wr S_{{\underline{\sigma}}(U)}\right)\cong\prod_{U\leq G}\left(U\wr S_{{\underline{\sigma}}(U)_{1}}\times U\wr S_{{\underline{\sigma}}(U)_{2}}\times\cdots\right)$
where $S_{{\underline{\sigma}}(U)}$ denotes the Young subgroup associated with the partition ${\underline{\sigma}}(U)$. By Lemma 2.1, a subset $Y$ of $W$ is ${\underline{\sigma}}$-transitive if and only if $Y$ is transitive on the cosets $W/H_{{\underline{\sigma}}}$. Thus, we do not get more notions of transitivity by considering compositions instead of partitions in the definition of the set $\Sigma_{n}(G)$ since the subgroups $H_{{\underline{\sigma}}}$ involved are conjugate to each other. Note that this also justifies that in Theorem 1.3 we only consider $(\sigma,k)$-flags where $\sigma$ is a partition, since the stabiliser of a $(\sigma,k)$-flag is $S_{\sigma}\times C_{2}\wr S_{k}$ where $S_{\sigma}=S_{\sigma_{1}}\times S_{\sigma_{2}}\times\cdots$ is a Young subgroup.
If $G$ is the group of one element, then ${\underline{\sigma}}\in\Sigma_{n}(G)$ is given by a partition $\sigma$ of $n$ via ${\underline{\sigma}}(G)=\sigma$ and the notion of a ${\underline{\sigma}}$-transitive set coincides with that of a $\sigma$-transitive subset of $S_{n}$ studied in [19].
We also emphasise the important special case that arises for $\lvert G\rvert>1$ when ${\underline{\sigma}}\in\Sigma_{n}(G)$ is given by ${\underline{\sigma}}(\{0\})=(1^{t})$ and ${\underline{\sigma}}(G)=(n-t)$. The corresponding subgroup is $G\wr S_{n-t}$ which appeared in Section 2.1. Then a ${\underline{\sigma}}$-transitive subset of $W$ is transitive on the set of $t$-tuples $((c_{1},i_{1}),\dots,(c_{t},i_{t}))$, where $i_{1},i_{2},\dots,i_{t}\in[n]$ are pairwise distinct integers and $c_{1},c_{2},\dots,c_{t}$ are elements of $G$. For $G=C_{r}$, we shall study these subsets in more detail in Section 6.
We can also give a geometric interpretation of some transitivity types using the $n$-dimensional cube.
Example 4.2.
Let $G=C_{2}$ and $n\in{\mathbb{N}}$ so that $W$ is the symmetry group of the $n$-dimensional cube. Consider ${\underline{\sigma}}\in\Sigma_{n}(G)$ given by ${\underline{\sigma}}(\{0\})=(t)$ and ${\underline{\sigma}}(C_{2})=(n-t)$. Then the stabiliser of a ${\underline{\sigma}}$-tabloid is the subgroup
$H_{t}=S_{t}\times C_{2}\wr S_{n-t}.$
The group $H_{t}$ is also the stabiliser of an $(n-t)$-dimensional face of the cube. Thus, a subset $Y$ of $W$ is ${\underline{\sigma}}$-transitive if and only if $Y$ is transitive on the $(n-t)$-dimensional faces of the $n$-dimensional cube.
We can also give an interpretation for the hyperoctahedron. If we view $C_{2}\wr S_{n}$ as the symmetry group of the hyperoctahedron, then the subgroup
$H_{t}=S_{t}\times C_{2}\wr S_{n-t}$
is the stabiliser of a $(t-1)$-dimensional face of the hyperoctahedron. Thus, a subset $Y$ of $W$ is ${\underline{\sigma}}$-transitive if and only if it is transitive on the $(t-1)$-dimensional faces of the hyperoctahedron.
In the next section, we show how these structures can be characterised as designs in an association scheme.
4.2. Decomposition of the permutation characters
For decomposing the permutation characters of the subgroups $H_{{\underline{\sigma}}}$, a certain algorithmically defined relation is helpful. A similar approach for decomposing permutation characters of the hyperoctahedral group $B_{n}$ (the case $G=C_{2}$) appears in [20], stated in a very restricted manner. We adapt this approach to wreath products and generalise it to cover all finite abelian groups $G$.
For a subgroup $U$ of $G$, we write
$U^{\circ}=\{\text{$g\in G:\theta^{g}(u)=1$ for all $u\in U$}\},$
so $U^{\circ}$ consists of the elements of $G$ that correspond to irreducible characters that are trivial on $U$. Now we can define a relation on $\Sigma_{n}(G)\times\Lambda_{n}(G)$.
Definition 4.3.
For ${\underline{\sigma}}\in\Sigma_{n}(G)$ and ${\underline{\lambda}}\in\Lambda_{n}(G)$, we write
${\underline{\sigma}}\;\rightarrow\;{\underline{\lambda}}$
if there exists ${\underline{\mu}}\in\Lambda_{n}(G)$ such that ${\underline{\mu}}(g)\unlhd{\underline{\lambda}}(g)$ for each $g\in G$ and ${\underline{\mu}}$ can be obtained from ${\underline{\sigma}}$ as follows.
  1. (1)
    For each $U\leq G$, colour each box of the Young diagram of ${\underline{\sigma}}(U)$ with an element of $U^{\circ}$.
  2. (2)
    The element ${\underline{\mu}}\in\Lambda_{n}(G)$ arises from the colouring from step (1) in the following way:
    For each $g\in G$, consider only the boxes coloured with $g$. Each row of each partition ${\underline{\sigma}}(U)$ with exactly $k$ boxes coloured with $g$ becomes a part of size $k$ in ${\underline{\mu}}(g)$ (the actual position of the elements in a row does not matter, only how often $g$ was placed there). In other words, ${\underline{\mu}}(g)$ has a part of size $k$ for every row of a Young diagram ${\underline{\sigma}}(U)$ with exactly $k$ appearances of $g$.
It is important to note that $\{0\}^{\circ}=G$ and $G^{\circ}=\{0\}$. This means that the boxes in ${\underline{\sigma}}(\{0\})$ can be coloured with every element of $G$ while the boxes in ${\underline{\sigma}}(G)$ all have to be coloured with the element 00.
Example 4.4.
We take $G=C_{6}$ and identify $C_{6}$ with $\{0,1,2,3,4,5\}$ in the natural way. Then the subgroups of $C_{6}$ are $\{0\}$, $\{0,3\}$, $\{0,2,4\}$, and $\{0,1,2,3,4,5\}$. We get
$\{0\}^{\circ}$ =$=\{0,1,2,3,4,5\},$
$\{0,3\}^{\circ}$ =$=\{0,2,4\},$
$\{0,2,4\}^{\circ}$ =$=\{0,3\},$
$\{0,1,2,3,4,5\}^{\circ}$ =$=\{0\}.$
These are the elements we can use in the colouring of the Young diagrams in step (1) of Definition 4.3. Consider ${\underline{\sigma}}\in\Sigma_{14}(C_{6})$ given by
$\{0\}$$\{0,3\}$$\{0,2,4\}$$\{0,1,2,3,4,5\}$
{ytableau}
{ytableau}
{ytableau}
{ytableau}
. A possible colouring of ${\underline{\sigma}}$ is given by
$\{0\}$$\{0,3\}$$\{0,2,4\}$$\{0,1,2,3,4,5\}$
{ytableau} 2 5
5 {ytableau} 4 2 2
4 4 {ytableau} 3 0 3
{ytableau} 0 0
0
. Then the corresponding ${\underline{\mu}}\in\Lambda_{14}(C_{6})$ is given by
00 $1$ $2$ $3$ $4$ $5$
[…]
This shows that ${\underline{\sigma}}\to{\underline{\mu}}$ where
${\underline{\mu}}(0)=(211),\;{\underline{\mu}}(1)=\emptyset,\;{\underline{\mu}}(2)=(21),\;{\underline{\mu}}(3)=(2),\;{\underline{\mu}}(4)=(21),\;{\underline{\mu}}(5)=(11).$
[…]
The reason for introducing the relation $\rightarrow$ is the following theorem. It says that the relation $\to$ precisely characterises the irreducible characters $\chi^{{\underline{\lambda}}}$ that appear in the decomposition of the permutation character $\xi_{H_{{\underline{\sigma}}}}$.
Theorem 4.5.
For each ${\underline{\sigma}}\in\Sigma_{n}(G)$ and each ${\underline{\lambda}}\in\Lambda_{n}(G)$, we have
$\big{\langle}\xi_{H_{\underline{\sigma}}},\chi^{{\underline{\lambda}}}\big{\rangle}\neq 0\;\Longleftrightarrow\;{\underline{\sigma}}\;\rightarrow\;{\underline{\lambda}}.$
Proof.
Let ${\underline{\lambda}}\in\Lambda_{n}(G)$. We pick a tuple $(g_{1},\ldots,g_{n})\in G^{n}$ such that every $g\in G$ occurs precisely $|{\underline{\lambda}}(g)|$ times. This defines a Young subgroup
$S=\prod_{g\in G}S_{g},$
where $S_{g}$ permutes the indices $i\in[n]$ with $g_{i}=g$. Hence, we have an ordered set partition $P$ of $[n]$ indexed by $G$. The part of $P$ indexed by $g\in G$ is the set of indices $i\in[n]$ with $g_{i}=g$.
We write $K=G\wr S=G^{n}\rtimes S$. By Theorem 3.1, we have $\chi^{\underline{\lambda}}=(\phi\cdot\rho)\uparrow_{K}^{W}$ where $\phi$ is the irreducible character of $G^{n}$ of type $(g_{1},\ldots,g_{n})$ and $\rho$ is the irreducible character of $S$ given by
$\rho=\bigotimes_{g\in G}\chi^{{\underline{\lambda}}(g)}.$
Write $H=H_{\underline{\sigma}}$. By Frobenius reciprocity, we have
$\big{\langle}\xi_{H},\chi^{{\underline{\lambda}}}\big{\rangle}=\big{\langle}1\uparrow_{H}^{W},\chi^{\underline{\lambda}}\big{\rangle}=\big{\langle}1\uparrow_{H}^{W},(\phi\cdot\rho)\uparrow_{K}^{W}\big{\rangle}=\big{\langle}1\uparrow_{H}^{W}\downarrow_{K}^{W},\phi\cdot\rho\big{\rangle}.$
For $s\in W$, write $H^{s}=sHs^{-1}$. Let $R$ be a complete set of double coset representatives of $K\,\backslash W/H$. Then by Mackey’s formula (Theorem 2.2)
$1\uparrow_{H}^{W}\downarrow_{K}^{W}=\sum_{s\in R}1\downarrow_{H^{s}\cap K}^{H^{s}}\uparrow_{H^{s}\cap K}^{K}=\sum_{s\in R}1\uparrow_{H^{s}\cap K}^{K}$
and Frobenius reciprocity, we have
(12) $\big{\langle}\xi_{H},\chi^{{\underline{\lambda}}}\big{\rangle}=\sum_{s\in R}\big{\langle}1\uparrow_{H^{s}\cap K}^{K},\phi\cdot\rho\big{\rangle}=\sum_{s\in R}\big{\langle}1,(\phi\cdot\rho)\downarrow_{H^{s}\cap K}^{K}\big{\rangle}.$
Note that all summands on the right hand side are nonnegative, so that the total sum is nonzero if and only if one summand is nonzero. Since $K=G\wr S$, we have $G^{n}\subseteq K$, so we may assume that all representatives in $R$ are elements of $S_{n}$. Hence, (12) is nonzero if and only if there exists $s\in S_{n}$ such that
$\big{\langle}1,(\phi\cdot\rho)\downarrow_{H^{s}\cap K}^{K}\big{\rangle}\neq 0,$
or equivalently
(13) $\big{\langle}1,\phi\downarrow_{H^{s}\cap G^{n}}^{G^{n}}\big{\rangle}\cdot\big{\langle}1,\rho\downarrow_{H^{s}\cap S}^{S}\big{\rangle}\neq 0.$
The left hand side of (13) is nonzero if and only if both factors are nonzero. We focus on the second factor first.
Note that $H\cap S_{n}$ is a Young subgroup of $S_{n}$ whose factors correspond to the parts of ${\underline{\sigma}}(U)$ as $U$ ranges over the subgroups of $G$. Since $s\in S_{n}$, the same is true for $(H\cap S_{n})^{s}=H^{s}\cap S_{n}$. Thus, there is an ordered set partition $Q$ of $[n]$ corresponding to the Young subgroup $H^{s}\cap S_{n}$ where every part is associated with a row of a Young diagram ${\underline{\sigma}}(U)$ for some subgroup $U$. Recall that $S$ is a Young subgroup with corresponding ordered set partition $P$ where every part is associated with an element $g\in G$. It follows that $H^{s}\cap S$ is a Young subgroup where the corresponding ordered set partition consists precisely of the parts $p\cap q$ with $p\in P$ and $q\in Q$. This ordered set partition defines a colouring of ${\underline{\sigma}}$ in the following way. For every element $p\cap q$ with $p\in P$ and $q\in Q$, we colour the boxes corresponding to $p\cap q$ in the row associated with $q$ with the element $g$ associated with $p$. Hence, we get ${\underline{\mu}}\in\Lambda_{n}(G)$ with $\lvert{\underline{\mu}}(g)\rvert=|{\underline{\lambda}}(g)|$ for all $g\in G$ and
(14) $H^{s}\cap S=\prod_{g\in G}S_{{\underline{\mu}}(g)}.$
In other words, ${\underline{\mu}}$ describes how the factors $S_{g}$ of $S$ decompose after intersecting $H^{s}$ with $S$. The parts of ${\underline{\mu}}(g)$ correspond to the number of times $g$ appears in a row of ${\underline{\sigma}}(U)$ for some $U\leq G$ in this colouring. Moreover, we have
$\big{\langle}1,\rho\downarrow_{H^{s}\cap S}^{S}\big{\rangle}$ $=\big{\langle}1,\bigotimes_{g\in G}\chi^{{\underline{\lambda}}(g)}\downarrow_{H^{s}\cap S}^{S}\big{\rangle}$
$=\prod_{g\in G}\big{\langle}1,\chi^{{\underline{\lambda}}(g)}\downarrow^{S_{g}}_{S_{{\underline{\mu}}(g)}}\big{\rangle}$
$=\prod_{g\in G}\big{\langle}1\uparrow^{S_{g}}_{S_{{\underline{\mu}}(g)}},\chi^{{\underline{\lambda}}(g)}\big{\rangle}.$
The left hand side is nonzero if and only if every factor on the right hand side is nonzero. By Young’s rule (see for example [21, Theorem 2.11.2]), this happens if and only if ${\underline{\mu}}(g)\unlhd{\underline{\lambda}}(g)$ for each $g\in G$.
Now we consider the first factor on the left hand side of (13). Since $\phi$ is an irreducible character of $G^{n}$ and therefore of degree $1$, it follows that the factor is nonzero if and only if $\phi$ is trivial on $H^{s}\cap G^{n}$. Since $H^{s}$ is conjugate to $H$, it is the stabiliser of a ${\underline{\sigma}}$-tabloid. Hence, $H^{s}\cap G^{n}$ is a direct product of subgroups of $G$, where $U\leq G$ occurs $\lvert{\underline{\sigma}}(U)\rvert$ times. This means that to each $i\in\{1,2,\dots,n\}$ we assign a subgroup $U$ of $G$, so that $H^{s}\cap G^{n}=U_{1}\times\cdots\times U_{n}$. Remember that the character $\phi$ is of type $(g_{1},\dots,g_{n})$. Since $\phi$ is trivial on $U_{1}\times\cdots\times U_{n}$, we find that $(g_{1},\dots,g_{n})\in U_{1}^{\circ}\times\cdots\times U_{n}^{\circ}$.
Now suppose that $\big{\langle}\xi_{H},\chi^{{\underline{\lambda}}}\big{\rangle}\neq 0$. Then both factors in (13) are nonzero. From (14), we obtain a colouring ${\underline{\mu}}\in\Lambda_{n}(G)$ of ${\underline{\sigma}}$. Since the first factor of (13) is nonzero, all rows of ${\underline{\sigma}}(U)$ are coloured with elements of $U^{\circ}$, corresponding to step (1) in Definition 4.3. Since the second factor of (13) is nonzero, we have ${\underline{\mu}}(g)\unlhd{\underline{\lambda}}(g)$ for every $g\in G$, corresponding to step (2) in Definition 4.3. Thus, we obtain ${\underline{\sigma}}\to{\underline{\lambda}}$.
For the reverse direction, assume that ${\underline{\sigma}}\to{\underline{\lambda}}$. Then there exists a colouring ${\underline{\mu}}\in\Lambda_{n}(G)$ of ${\underline{\sigma}}$ obtained from steps (1) and (2) in Definition 4.3. Remember that $H\cap S_{n}$ is a Young subgroup of $S_{n}$ whose parts correspond to the parts of ${\underline{\sigma}}(U)$ for $U\leq G$. Now consider the Young subgroup $T=\prod_{g\in G}T_{g}$ of $S_{n}$ where each part $T_{g}$ permutes the boxes of ${\underline{\sigma}}$ coloured with $g\in G$. By the same reasoning as for the subgroup $H^{s}\cap S$, we find that $H\cap T$ is a Young subgroup of the form $\prod_{g\in G}S_{{\underline{\mu}}(g)}$. Hence, it is conjugate to the subgroup in (14) and there exists $s\in S_{n}$ such that
$(H\cap T)^{s}=H^{s}\cap T^{s}=\prod_{g\in G}S_{{\underline{\mu}}(g)}$
is the subgroup from (14). Note that $T^{s}$ is conjugate to $S$, so we can replace $S$ by $T^{s}$ in the beginning of the proof and do the same calculations for $T^{s}$ as for $S$. The character $\rho$ is the same, while the type of the character $\phi$ is a permutation of the type $(g_{1},\ldots,g_{n})$. In view of Theorem 3.1, this does not change the character $\chi^{{\underline{\lambda}}}$. Thus, we find that equation (14) holds. Since ${\underline{\mu}}(g)\unlhd{\underline{\lambda}}(g)$ for each $g\in G$, the second factor of (13) is nonzero. Moreover, $\phi$ is trivial on $H^{s}\cap G^{n}$, so the first factor is nonzero. All in all, we obtain $\big{\langle}\xi_{H},\chi^{{\underline{\lambda}}}\big{\rangle}\neq 0$. ∎
We can use Theorem 4.5 to obtain a characterisation of ${\underline{\sigma}}$-transitive sets in $W$.
Theorem 4.6.
Let ${\underline{\sigma}}\in\Sigma_{n}(G)$ and let $Y$ be a nonempty subset of $W$ with dual distribution $(a^{\prime}_{\underline{\lambda}})$. Then $Y$ is ${\underline{\sigma}}$-transitive if and only if
$a^{\prime}_{\underline{\lambda}}=0\quad\text{for all ${\underline{\lambda}}\in\Lambda_{n}(G)$ satisfying ${\underline{\sigma}}\to{\underline{\lambda}}$ and ${\underline{\lambda}}(0)\neq(n)$},$
where 00 is the identity of $G$.
Proof.
Follows immediately from Theorem 2.5 and Theorem 4.5. ∎
[…]
Theorem 4.6 is crucial in order to derive a generalisation of the Livingstone-Wagner theorem in Section 5 and to describe $t$-transitive sets in $C_{r}\wr S_{n}$ in Section 6.
4.3. Cliques
In this section we consider so-called cliques in $W$ and discuss their relationship to transitive subsets.
Definition 4.7.
Let ${\underline{\sigma}}\in\Sigma_{n}(G)$. A non-empty subset $Y$ of $W$ is called a ${\underline{\sigma}}$-clique if for every pair of distinct elements $x,y\in Y$ there is no ${\underline{\sigma}}$-tabloid that is fixed by $x^{-1}y$.
We are interested in an algebraic characterisation of ${\underline{\sigma}}$-cliques as cliques in an association scheme. This comes down to determining the conjugacy classes of the elements in the stabiliser $H_{{\underline{\sigma}}}$ of a ${\underline{\sigma}}$-tabloid. Note that all stabilisers of ${\underline{\sigma}}$-tabloids are conjugate to each other, so the answer does not depend on the specific ${\underline{\sigma}}$-tabloid chosen. Remember that $H_{{\underline{\sigma}}}$ is an inner direct product of subgroups of the form $U\wr S_{k}$ for a subgroup $U\leq G$ and a positive integer $k$. The conjugacy classes of the group $U\wr S_{k}$ are indexed by the set $\Lambda_{k}(U)$. We interpret an element ${\underline{\lambda}}\in\Lambda_{k}(U)$ as an element of $\Lambda_{k}(G)$ by defining ${\underline{\lambda}}(g)=\emptyset$ for every $g\in G$ with $g\not\in U$. Moreover, for $k<n$ we interpret an element ${\underline{\lambda}}\in\Lambda_{k}(G)$ as an element $\widetilde{{\underline{\lambda}}}\in\Lambda_{n}(G)$ by defining $\widetilde{{\underline{\lambda}}}(0)={\underline{\lambda}}(0)\cup 1^{n-k}$ and $\widetilde{{\underline{\lambda}}}(g)={\underline{\lambda}}(g)$ for all $g\in G\,\setminus\{0\}$.
For two elements ${\underline{\lambda}},{\underline{\mu}}\in\Lambda_{n}(G)$, we define the element ${\underline{\lambda}}\cup{\underline{\mu}}\in\Lambda_{m+n}(G)$ via $\left({\underline{\lambda}}\cup{\underline{\mu}}\right)(g)={\underline{\lambda}}(g)\cup{\underline{\mu}}(g)$ for every $g\in G$. We can now describe the conjugacy classes of the direct product of two groups of the form $U\wr S_{k}$.
Lemma 4.8.
Let $U,V\leq G$ be subgroups of $G$ and $k,\ell\in{\mathbb{N}}$ with $k+\ell\leq n$. Consider the subgroup $H=\left(U\wr S_{A}\right)\times\left(V\wr S_{B}\right)$ of $G\wr S_{n}$ where $A,B\subseteq[n]$ are disjoint subsets with $|A|=k$ and $|B|=\ell$. Then the conjugacy classes of $H$ are indexed by the set
$\left\{{\underline{\lambda}}\cup{\underline{\mu}}:{\underline{\lambda}}\in\Lambda_{k}(U),\;{\underline{\mu}}\in\Lambda_{\ell}(V)\right\}.$
Proof.
Let $x=(f,\pi)\in U\wr S_{A}$ and $y=(g,\sigma)\in V\wr S_{B}$. Then their product is given by
$xy=(f+g^{\pi},\pi\sigma).$
Here, we have $f=(f_{1},\ldots,f_{n})\in G^{n}$ with $f_{a}\in U$ for $a\in A$ and $f_{i}=0$ otherwise. Similarly, we have $g=(g_{1},\ldots,g_{n})\in G^{n}$ with $g_{b}\in V$ for $b\in B$ and $g_{i}=0$ otherwise. Note that since $A$ and $B$ are disjoint, $\pi$ acts trivially on $g$ and we find that
$xy=(f+g,\pi\sigma).$
Recall that the conjugacy classes of $x$ and $y$ are given by the cycle lengths of $\pi$ and $\sigma$, respectively, together with the signs of the cycles (see Section 3.2). Since $S_{A}$ and $S_{B}$ permute disjoint subsets of $[n]$, we find that the set of cycle lengths of $\pi\sigma$ is simply the disjoint union of the set of cycle lengths of $\pi$ and the set of cycle lengths of $\sigma$. Moreover, since $A$ and $B$ are disjoint, the cycle signs of $\pi$ and $\sigma$ do not change. Hence, if $x$ is of cycle type ${\underline{\lambda}}$ and $y$ is of cycle type ${\underline{\mu}}$, then $xy$ is of cycle type ${\underline{\lambda}}\cup{\underline{\mu}}$. The cycle type of $xy$ as an element of $G\wr S_{n}$ is then obtained by adding cycles of length 1 for all elements of $[n]\,\setminus\left(A\cup B\right)$ (if $k+\ell<n$). This corresponds exactly to how we identified elements of $\Lambda_{k+\ell}(G)$ with elements of $\Lambda_{n}(G)$, proving the theorem. ∎
Denote by $C({\underline{\sigma}})$ the set of all ${\underline{\lambda}}\in\Lambda_{n}(G)$ such that there exists an element in $H_{{\underline{\sigma}}}$ that lies in the conjugacy class $C_{{\underline{\lambda}}}$ of $W$. Applying Lemma 4.8 repeatedly immediately gives the following result.
Theorem 4.9.
Let ${\underline{\sigma}}\in\Lambda_{n}(G)$. Consider a stabiliser $H_{{\underline{\sigma}}}$ of a ${\underline{\sigma}}$-tabloid and write
$H_{{\underline{\sigma}}}=H_{1}\times\dots\times H_{\ell}$
for some $\ell\in{\mathbb{N}}$ where $H_{i}=U_{i}\wr S_{k_{i}}$ for some subgroup $U_{i}\leq G$ and $k_{i}\in{\mathbb{N}}$. Then we have
$C({\underline{\sigma}})=\left\{\bigcup_{i=1}^{\ell}{\underline{\lambda}}^{i}:{\underline{\lambda}}^{i}\in\Lambda_{k_{i}}(U_{i})\right\}.$
The following result should be compared to Theorem 4.6 and highlights that the concept of a ${\underline{\sigma}}$-clique is dual to the concept of ${\underline{\sigma}}$-transitivity. Recall that the conjugacy class of the identity of $W$ is indexed by the unique element ${\underline{\lambda}}\in\Lambda_{n}(G)$ with ${\underline{\lambda}}(0)=(1^{n})$.
Theorem 4.10.
Let ${\underline{\sigma}}\in\Sigma_{n}(G)$ and let $Y$ be a non-empty subset of $W$ with inner distribution $(a_{{\underline{\lambda}}})$. Then $Y$ is a ${\underline{\sigma}}$-clique if and only if
$a_{\underline{\lambda}}=0\quad\text{for all ${\underline{\lambda}}\in C({\underline{\sigma}})$ with ${\underline{\lambda}}(0)\neq(1^{n})$}.$
Proof.
Let $x,y\in Y$ be distinct elements of $Y$. Then the conjugacy class of $x^{-1}y$ is not that of the identity of $W$. If $a_{{\underline{\lambda}}}=0$ for every ${\underline{\lambda}}\in C({\underline{\sigma}})$ with ${\underline{\lambda}}(0)\neq(1^{n})$, then the conjugacy class of $x^{-1}y$ is not indexed by an element ${\underline{\lambda}}\in C({\underline{\sigma}})$ because in view of (5), $x^{-1}y\in C_{{\underline{\lambda}}}$ implies $a_{{\underline{\lambda}}}>0$. It follows that $x^{-1}y$ cannot be an element of $H^{w}=wH_{{\underline{\sigma}}}w^{-1}$ for any $w\in W$. Since the groups $H^{w}$ for $w\in W$ are exactly all the stabilisers of ${\underline{\sigma}}$-tabloids, $x^{-1}y$ does not fix a ${\underline{\sigma}}$-tabloid and thus $Y$ is a ${\underline{\sigma}}$-clique.
For the converse direction, note that if $x\in C_{{\underline{\lambda}}}$ stabilises a ${\underline{\sigma}}$-tabloid $T$, then $wxw^{-1}$ stabilises the ${\underline{\sigma}}$-tabloid $wT$. Hence, either all elements of a conjugacy class $C_{{\underline{\lambda}}}$ stabilise a ${\underline{\sigma}}$-tabloid or none. Thus, if $Y$ is a ${\underline{\sigma}}$-clique, then the conjugacy class of the quotient $x^{-1}y$ for distinct $x,y\in Y$ cannot be a conjugacy class of an element of $H_{{\underline{\sigma}}}$. This shows that $a_{{\underline{\lambda}}}=0$ for ${\underline{\lambda}}\in C({\underline{\sigma}})$ with ${\underline{\lambda}}(0)\neq(1^{n})$. ∎
We now establish a connection between ${\underline{\sigma}}$-transitive sets and ${\underline{\sigma}}$-cliques in $W$.
Theorem 4.11.
Let ${\underline{\sigma}}\in\Lambda_{n}(G)$, let $H_{{\underline{\sigma}}}$ be the stabiliser of a ${\underline{\sigma}}$-tabloid and let $Y$ be a non-empty subset of $W$.
  1. (a)
    If $Y$ is a ${\underline{\sigma}}$-clique, then $|Y|\leq|W|/|H|$ with equality if and only if $Y$ is ${\underline{\sigma}}$-transitive.
  2. (b)
    If $Y$ is ${\underline{\sigma}}$-transitive, then $|Y|\geq|W|/|H|$ with equality if and only if $Y$ is a ${\underline{\sigma}}$-clique.
In both cases, equality implies that $Y$ is sharply ${\underline{\sigma}}$-transitive.
Proof.
Observe that for each $(x,y)\in H\times Y$ there is precisely one element $w\in W$ with $wx=y$ (namely $w=yx^{-1}$). Thus, we have
(15) $\sum_{w\in W}\left|Y\cap wH\right|=|Y|\cdot|H|.$
The quotient of any two elements in $Y\cap wH$ fixes a ${\underline{\sigma}}$-tabloid. Hence, if $Y$ is a ${\underline{\sigma}}$-clique, then every summand on the left hand side of (15) is at most 1. This gives $|Y|\cdot|H|\leq|W|$, proving the bound in (a). Now, since $H$ is the stabiliser of a ${\underline{\sigma}}$-tabloid $T$, the coset $wH$ contains precisely the elements of $W$ that map $T$ to $wT$. Thus, if $Y$ is ${\underline{\sigma}}$-transitive, then every summand on the left hand side of (15) is at least 1. This gives $|Y|\cdot|H|\geq|W|$, proving the bound in (b).
In both cases, equality occurs if and only if $|Y\cap w\widetilde{H}|=1$ for every $w\in W$ and every stabiliser $\widetilde{H}$ of a ${\underline{\sigma}}$-tabloid. This is equivalent to $Y$ being sharply ${\underline{\sigma}}$-transitive. ∎
Theorem 4.11 can also be obtained using the so-called clique-coclique bound (see [8, Theorem 3.9]).
5. Comparison of transitivity types
We now turn to generalisations of Theorem 1.1 by Livingstone and Wagner. Let ${\underline{\sigma}},{\underline{\tau}}\in\Sigma_{n}(G)$. We write ${\underline{\sigma}}\succeq{\underline{\tau}}$ if for every subset $Y$ of $W$ we have
[…]
Hence, writing
$M_{{\underline{\sigma}}}=\{{\underline{\lambda}}\in\Lambda_{n}(G):{\underline{\sigma}}\to{\underline{\lambda}}\},$
Theorem 4.6 implies
${\underline{\sigma}}\succeq{\underline{\tau}}\;\Longleftrightarrow\;M_{{\underline{\sigma}}}\supseteq M_{{\underline{\tau}}}.$
We shall give a characterisation of the relation $\succeq$ in special cases. We say that ${\underline{\sigma}}\in\Sigma_{n}(G)$ is simple if ${\underline{\sigma}}(U)=\emptyset$ for each nontrivial proper subgroup $U$ of $G$. If $|G|>1$, then the shape of a simple ${\underline{\sigma}}\in\Sigma_{n}(G)$ is the pair $(\sigma,\rho)$, where ${\underline{\sigma}}(\{0\})=\sigma$ and ${\underline{\sigma}}(G)=\rho$. In order to fit the case $|G|=1$ into this framework, we define the shape of a partition $\sigma\in\Sigma_{n}(\{0\})$ to be $(\sigma,\emptyset)$. We shall identify a simple ${\underline{\sigma}}\in\Sigma_{n}(G)$ with its shape. We define the size of a double partition $(\sigma,\rho)$ to be $\lvert\sigma\rvert+\lvert\rho\rvert$. For double partitions $(\sigma,\rho)$ and $(\tau,\pi)$ of the same size $n$, we write $(\sigma,\rho)\succeq(\tau,\pi)$ if their corresponding symbols in $\Sigma_{n}(G)$ are related in this way. Note that if $G$ has prime order, then simplicity of ${\underline{\sigma}}\in\Sigma_{n}(G)$ is no restriction.
We call ${\underline{\sigma}}\in\Sigma_{n}(G)$ parabolic if it is simple of shape $(\sigma,(k))$ for a non-negative integer $k$ (the case $k=0$ should be read as $(\sigma,\emptyset)$). We write $(\sigma,k)$ for the shape $(\sigma,(k))$. The term ”parabolic” stems from the fact that if ${\underline{\sigma}}$ is parabolic in the hyperoctahedral group, the case $G=C_{2}$, then the corresponding stabiliser $H_{{\underline{\sigma}}}$ is a parabolic subgroup if $W$ is viewed as a Coxeter group. It turns out that parabolic ${\underline{\sigma}}$ are especially nice to work with and capture important transitivity types like $t$-transitivity for ${\underline{\sigma}}=((1^{t}),n-t)$ and $t$-homogeneity for ${\underline{\sigma}}=((t),n-t)$.
We start with some simple yet helpful properties of the relation $\succeq$.
Proposition 5.1.
Let $(\sigma,\rho)$ and $(\tau,\pi)$ be two double partitions of the same size satisfying $(\sigma,\rho)\succeq(\tau,\pi)$. Then we have
  1. (a)
    $\rho\unlhd\pi$,
  2. (b)
    $\sigma^{\prime}\unrhd\tau^{\prime}$,
  3. (c)
    $\sigma\cup\rho\unlhd\tau\cup\pi$.
Proof.
To prove (a) and (b), consider ${\underline{\lambda}}\in\Lambda_{n}(G)$ given by ${\underline{\lambda}}(0)=\pi$ and ${\underline{\lambda}}(g)=\tau$ for some fixed nonzero $g\in G$ (in the case $G=\{0\}$, we have $\pi=\emptyset$ and use ${\underline{\lambda}}\in\Lambda_{n}(\{0\})$ with ${\underline{\lambda}}(0)=\tau$). Then we have $(\tau,\pi)\to{\underline{\lambda}}$, and since $(\sigma,\rho)\succeq(\tau,\pi)$, we must also have $(\sigma,\rho)\to{\underline{\lambda}}$. Let ${\underline{\mu}}\in\Lambda_{n}(G)$ be obtained from any colouring of $(\sigma,\rho)$. Then the parts of $\rho$ must occur as parts in ${\underline{\mu}}(0)$. Hence $\rho\unlhd{\underline{\lambda}}(0)$ and so $\rho\unlhd\pi$, proving (a). On the other hand, for every $k$ the number of boxes in the first $k$ columns of ${\underline{\mu}}(g)$ cannot be larger than the number of boxes in the first $k$ columns of $\sigma$. Hence $\sigma^{\prime}\unrhd{\underline{\lambda}}(g)^{\prime}$ and so $\sigma^{\prime}\unrhd\tau^{\prime}$, proving (b).
To prove (c), consider ${\underline{\lambda}}\in\Lambda_{n}(G)$ given by ${\underline{\lambda}}(0)=\tau\cup\pi$. Then we have $(\tau,\pi)\to{\underline{\lambda}}$ and so $(\sigma,\rho)\to{\underline{\lambda}}$. This forces $\sigma\cup\rho\unlhd\tau\cup\pi$. ∎
The following example shows that the conditions (a), (b), (c) in Proposition 5.1 are not sufficient to guarantee $(\sigma,\rho)\succeq(\tau,\pi)$. Take $(\tau,\pi)=((22),(311))$ and $(\sigma,\rho)=((311),(22))$. Let ${\underline{\lambda}}\in\Lambda_{9}(G)$ be given by ${\underline{\lambda}}(0)=(31^{4})$ and ${\underline{\lambda}}(g)=(1^{2})$ for some fixed nonzero $g\in G$. Then $(\tau,\pi)\to{\underline{\lambda}}$ holds, but $(\sigma,\rho)\to{\underline{\lambda}}$ does not hold. Hence $(\sigma,\rho)\succeq(\tau,\pi)$ is not true.
We shall however prove that the conditions (a) and (c) in Proposition 5.1 are sufficient to guarantee $(\sigma,\rho)\succeq(\tau,\pi)$ in the parabolic case.
Let us first investigate how the relation $\succeq$ behaves when changing the shape $(\sigma,\rho)$ slightly.
Lemma 5.2 (Dominance rule).
Let $(\sigma,\rho)$ and $(\tau,\pi)$ be two double partitions of the same size satisfying $\lvert\sigma\rvert=\lvert\tau\rvert$, $\sigma\unlhd\tau$ and $\rho\unlhd\pi$. Then we have $(\sigma,\rho)\succeq(\tau,\pi)$.
Proof.
Since $\succeq$ is a transitive relation, we can show $(\sigma,\rho)\succeq(\tau,\pi)$ in two steps via $(\sigma,\rho)\succeq(\sigma,\pi)\succeq(\tau,\pi)$.
First, we show that $(\sigma,\rho)\succeq(\sigma,\pi)$. Consider ${\underline{\lambda}}\in\Lambda_{n}(G)$ with $(\sigma,\pi)\to{\underline{\lambda}}$. We have to show $(\sigma,\rho)\to{\underline{\lambda}}$. According to Definition 4.3, there is ${\underline{\beta}}\in\Lambda_{n}(G)$ arising from a colouring $C$ of the boxes of $(\sigma,\pi)$ such that ${\underline{\beta}}(g)\unlhd{\underline{\lambda}}(g)$ for all $g\in G$. We need to find a colouring $\widetilde{C}$ of $(\sigma,\rho)$ in step (1) of Definition 4.3 such that we obtain $(\sigma,\rho)\to{\underline{\lambda}}$. Define $\widetilde{C}$ by colouring the boxes of $\sigma$ in $(\sigma,\rho)$ in the same way as in the colouring $C$ of $(\sigma,\pi)$. Colour all boxes of $\rho$ with 00 (this is forced because of $G^{\circ}=\{0\}$). This assignment defines a colouring $\widetilde{C}$ of $(\sigma,\rho)$ which gives $\underline{\gamma}\in\Lambda_{n}(G)$ with $\underline{\gamma}(g)={\underline{\beta}}(g)$ for $g\neq 0$. Moreover, writing $\alpha$ for the partition of boxes of $\sigma$ that are coloured with 0, we have $\underline{\gamma}(0)=\alpha\cup\rho$ and ${\underline{\beta}}(0)=\alpha\cup\pi$. Since we have $\rho\unlhd\pi$, we obtain $\underline{\gamma}(0)\unlhd{\underline{\beta}}(0)$. Hence, we find that $\underline{\gamma}(g)\unlhd{\underline{\lambda}}(g)$ for every $g\in G$, proving $(\sigma,\rho)\to{\underline{\lambda}}$.
Now we show $(\sigma,\pi)\succeq(\tau,\pi)$. Consider ${\underline{\lambda}}\in\Lambda_{n}(G)$ with $(\tau,\pi)\to{\underline{\lambda}}$. We have to show $(\sigma,\pi)\to{\underline{\lambda}}$. Since $\sigma\unlhd\tau$, the Young diagram of $\tau$ can be obtained from the Young diagram of $\sigma$ by moving up boxes one at a time. Since the relation $\succeq$ is transitive, it is enough to consider a single box that is moved up, say from row $j$ to row $i$. Thus, we can assume that $\tau_{i}=\sigma_{i}+1$, $\tau_{j}=\sigma_{j}-1$ and $\tau_{k}=\sigma_{k}$ for every $k\neq i,j$.
According to Definition 4.3, there is ${\underline{\beta}}\in\Lambda_{n}(G)$ arising from a colouring $C$ of the boxes of $(\tau,\pi)$ such that ${\underline{\beta}}(g)\unlhd{\underline{\lambda}}(g)$ for all $g\in G$. Consider the colouring of the boxes of $\tau$ in row $i$ and row $j$. Since $\sigma_{i}\geq\sigma_{j}$, we have $\tau_{i}\geq\tau_{j}+2$. Thus, there exists an element $g\in G$ such that in row $i$ of $\tau$ there are more boxes coloured with $g$ than in row $j$ of $\tau$.
Now define a colouring $\widetilde{C}$ of $(\sigma,\pi)$ in the following way. Colour the box that was moved to obtain $\tau$ from $\sigma$ with $g$ and colour all other boxes of $\sigma$ in the same way that they were coloured in $\tau$ in the colouring $C$. All boxes of $\pi$ have to be coloured with 00. This colouring $\widetilde{C}$ of $(\sigma,\pi)$ gives $\underline{\gamma}\in\Lambda_{n}(G)$ with $\underline{\gamma}(h)={\underline{\beta}}(h)$ for every $h\neq g$. Now consider ${\underline{\beta}}(g)$ and $\underline{\gamma}(g)$. Let $x$ denote the number of times that $g$ appears in row $j$ of $\tau$ in the colouring $C$ giving rise to ${\underline{\beta}}$. By the choice of $g$, we find that $g$ appears $x+k$ times in row $i$ of $\tau$ for some $k\geq 1$. Thus, ${\underline{\beta}}(g)$ contains the parts $x$ and $x+k$. By construction of $\widetilde{C}$, $\underline{\gamma}(g)$ contains the parts $x+1$ and $x+k-1$. Since all other parts of ${\underline{\beta}}(g)$ and $\underline{\gamma}(g)$ are the same, we get $\underline{\gamma}(g)\unlhd{\underline{\beta}}(g)$. Hence, we find that $\underline{\gamma}(g)\unlhd{\underline{\lambda}}(g)$ for every $g\in G$, proving $(\sigma,\pi)\to{\underline{\lambda}}$. ∎
Proposition 5.1 and Lemma 5.2 immediately give a characterisation of the relation $\succeq$ in the case $|\sigma|=|\tau|$.
Corollary 5.3.
Let $(\sigma,\rho)$ and $(\tau,\pi)$ be two double partitions of the same size. If $|\sigma|=|\tau|$ ((((thus $|\rho|=|\pi|)$, then
$(\sigma,\rho)\succeq(\tau,\pi)\;\Longleftrightarrow\;\sigma\unlhd\tau\;\text{and}\;\rho\unlhd\pi.$
If $\sigma$ is a partition and $k$ and $r$ are positive integers, we denote with $\sigma+_{r}k$ the partition obtained by adding $k$ to the part $\sigma_{r}$ of $\sigma$ and then possibly reordering the parts to obtain a partition. In other words, we add the sequence $0^{r-1}k$ to the sequence $\sigma$ and reorder $\sigma$ such that it becomes a partition.
The next lemma describes a way to move some boxes from the partition $\sigma$ to the partition $\rho$.
Lemma 5.4 (Sum rule).
Let $(\sigma,\rho)$ be a double partition satisfying $\sigma_{r}\leq\rho_{s}$ for some positive integers $r$ and $s$. Then, for each nonnegative integer $k$, we have
$(\sigma+_{r}k,\rho)\succeq(\sigma,\rho+_{s}k).$
Proof.
Let ${\underline{\lambda}}\in\Lambda_{n}(G)$ with $(\sigma,\rho+_{s}k)\to{\underline{\lambda}}$. We have to show that $(\sigma+_{r}k,\rho)\to{\underline{\lambda}}$. According to Definition 4.3, there is ${\underline{\beta}}\in\Lambda_{n}(G)$ arising from a colouring $C$ of $(\sigma,\rho+_{s}k)$ such that ${\underline{\beta}}(g)\unlhd{\underline{\lambda}}(g)$ for all $g\in G$.
Define a colouring $\widetilde{C}$ of $(\sigma+_{r}k,\rho)$ as follows. Colour all boxes of $\sigma+_{r}k$ that are also boxes of $\sigma$ in the same way they are coloured in the colouring $C$ of $(\sigma,\rho+_{s}k)$. Colour the $k$ remaining boxes (that is the boxes that were added to the part $\sigma_{r}$) and all boxes of $\rho$ with the element 0 (the colouring of the boxes of $\rho$ is forced). This colouring $\widetilde{C}$ of $(\sigma+_{r}k,\rho)$ gives $\underline{\gamma}\in\Lambda_{n}(G)$ with $\underline{\gamma}(g)={\underline{\beta}}(g)$ for all $g\neq 0$.
Note that in the part $\sigma_{r}+k$, there are $x+k$ boxes coloured with $0\in G$ for some integer $x$ satisfying $0\leq x\leq\sigma_{r}$. Hence, ${\underline{\beta}}(0)$ contains the parts $x$ and $\rho_{s}+k$ while $\underline{\gamma}(0)$ contains the parts $x+k$ and $\rho_{s}$. We now distinguish the order of these parts.
If $x+k\leq\rho_{s}$, then $(\rho_{s},x+k)$ is a partition satisfying $(\rho_{s},x+k)\unlhd(\rho_{s}+k,x)$. If $x+k>\rho_{s}$, then $(x+k,\rho_{s})$ is a partition, which satisfies $(x+k,\rho_{s})\unlhd(\rho_{s}+k,x)$ since $x\leq\sigma_{r}\leq\rho_{s}$. In both cases, we find that $\underline{\gamma}(0)\unlhd{\underline{\beta}}(0)$ and so $\underline{\gamma}(0)\unlhd{\underline{\lambda}}(0)$. Hence, we find that $\underline{\gamma}(g)\unlhd{\underline{\lambda}}(g)$ for every $g\in G$, proving $(\sigma+_{r}k,\rho)\to{\underline{\lambda}}$. ∎
We emphasise two special cases of the sum rule.
Corollary 5.5 (Union rule).
Let $(\sigma,\rho)$ be a double partition. Then for every partition $\alpha$ we have
$(\sigma\cup\alpha,\rho)\succeq(\sigma,\rho\cup\alpha).$
Proof.
Since the relation $\succeq$ is transitive, it is enough to consider $\alpha=(k)$ for some integer $k$. Take $r$ and $s$ in the sum rule (Lemma 5.4) such that $\sigma_{r}=\rho_{s}=0$. Then $\sigma+_{r}k=\sigma\cup(k)$ and $\rho+_{s}k=\rho\cup(k)$. ∎
Corollary 5.6 (Exchange rule).
Let $(\sigma,\rho)$ be a double partition and let $i,j$ be integers with $\sigma_{i}>\rho_{j}$. Define the double partition $(\widetilde{\sigma},\widetilde{\rho})$ as the double partition $(\sigma,\rho)$ but with the part $\sigma_{i}$ of $\sigma$ and the part $\rho_{j}$ of $\rho$ exchanged (then possibly reordering the parts to obtain a partition). Then we have $(\sigma,\rho)\succeq(\widetilde{\sigma},\widetilde{\rho})$.
Proof.
Write $\sigma_{i}=\rho_{j}+(\sigma_{i}-\rho_{j})$ and take $k=\sigma_{i}-\rho_{j}$ in the sum rule (Lemma 5.4). ∎
The next theorem characterises the relation $\succeq$ for parabolic types.
Theorem 5.7.
Let $(\sigma,k)$ and $(\tau,\ell)$ be two double partitions of the same size. Then
$(\sigma,k)\succeq(\tau,\ell)\;\Longleftrightarrow\;\text{$k\leq\ell$ and $(\sigma\cup(k))\unlhd(\tau\cup(\ell))$}.$
Proof.
The forward direction follows from Proposition 5.1.
For the reverse direction, assume that $k\leq\ell$ and $(\sigma\cup(k))\unlhd(\tau\cup(\ell))$. It follows that $\sigma^{\prime}+1^{k}\unrhd\tau^{\prime}+1^{\ell}$ and therefore
$\sigma^{\prime}\unrhd\tau^{\prime}+0^{k}1^{\ell-k}.$
Hence
$\sigma\unlhd\tau\cup(0^{k},\ell-k),$
where $\tau\cup(0^{k},\ell-k)$ denotes the partition obtained by adding $l-k$ to the biggest part of $\tau$ that is less than or equal to $k$.
Let $a$ be the smallest positive integer such that $\tau_{a}\leq\ell$ and $b$ be the smallest positive integer such that $\tau_{b}\leq k$. Then we have
$\tau\cup(0^{k},\ell-k)=\tau+_{a}(\ell-\tau_{a})+_{a+1}(\tau_{a}-\tau_{a+1})+_{a+2}\cdots+_{b-1}(\tau_{b-2}-\tau_{b-1})+_{b}(\tau_{b-1}-k).$
Note that by construction, no parts need to be reordered after the operations $+_{i}$, so the order of operations does not matter. Hence, by first applying the dominance rule (Lemma 5.2) and then the sum rule (Lemma 5.4) multiple times, we obtain
$(\sigma,k)$ $\succeq\big{(}\tau+_{a}(\ell-\tau_{a})+_{a+1}(\tau_{a}-\tau_{a+1})+_{a+2}\cdots+_{b-1}(\tau_{b-2}-\tau_{b-1})+_{b}(\tau_{b-1}-k),k\big{)}$
$\succeq\big{(}\tau+_{a}(\ell-\tau_{a})+_{a+1}(\tau_{a}-\tau_{a+1})+_{a+2}\cdots+_{b-1}(\tau_{b-2}-\tau_{b-1}),\tau_{b-1}\big{)}\quad\text{as $\tau_{b}\leq k$}$
$\succeq\big{(}\tau+_{a}(\ell-\tau_{a})+_{a+1}(\tau_{a}-\tau_{a+1})+_{a+2}\cdots,\tau_{b-2}\big{)}\quad\text{as $\tau_{b-1}\leq\tau_{b-1}$}$
[…]
$\succeq\big{(}\tau+_{a}(\ell-\tau_{a}),\tau_{a}\big{)}\quad\text{since $\tau_{a+1}\leq\tau_{a+1}$}$
$\succeq\big{(}\tau,\ell\big{)}\quad\text{as $\tau_{a}\leq\tau_{a}$},$
as required. ∎
This theorem gives an easy criterion to compare parabolic transitivity types in the group $W$. It is a generalisation of the Livingstone-Wagner theorem (see Theorem 1.1) to the group $W$. Moreover, in the case $G=C_{1}$, the parabolic types become the double partitions $(\sigma,\emptyset)$, so we have $k=0$ and Theorem 4.10 specialises to Theorem 5 in [19], the generalisation of the Livingstone-Wagner theorem by Martin and Sagan for the symmetric group $S_{n}$.
Let us apply Theorem 5.7 to $W=C_{2}\wr S_{n}$, the symmetry group of the $n$-dimensional cube.
Example 5.8.
Let $W=C_{2}\wr S_{n}$ be the hyperoctahedral group, also denoted as $B_{n}$. Then all types ${\underline{\sigma}}$ are simple.
First, let $n=3$ and consider the parabolic types ${\underline{\sigma}}=(2,1)$ and ${\underline{\tau}}=(1,2)$. They correspond to the subgroups $H_{{\underline{\sigma}}}=S_{2}\times B_{1}$ and $H_{{\underline{\tau}}}=S_{1}\times B_{2}$. Geometrically, $H_{{\underline{\sigma}}}$ is the stabiliser of an edge of the cube and $H_{{\underline{\tau}}}$ is the stabiliser of a face of the cube. Thus, the double partition $(2,1)$ stands for edge transitivity and the double partition $(1,2)$ stands for face transitivity. Applying Theorem 5.7 gives $(2,1)\succeq(1,2)$, that is every $(2,1)$-transitive set is also $(1,2)$-transitive. Geometrically, this means that every subset of $C_{2}\wr S_{3}$ that acts transitively on the edges of a 3-dimensional cube also acts transitively on the faces of a 3-dimensional cube.
We can generalise this observation to the $n$-dimensional cube. The stabiliser of a $t$-dimensional face of the cube is $S_{n-t}\times B_{t}$. Thus, the double partition $(n-t,t)$ describes transitivity on the $t$-dimensional faces of the $n$-dimensional cube. Theorem 5.7 gives
$(n-t,t)\succeq(n-s,s)\;\Longleftrightarrow\;t\leq s\text{ and }((n-t)\cup(t))\unlhd((n-s)\cup(s)).$
We distinguish the cases $t\geq n-t$ and $t<n-t$.
If $t\geq n-t$, then Theorem 5.7 gives $(n-t,t)\succeq(n-s,s)$ for every $s\geq t$ and $(n-t,t)\not\succeq(n-s,s)$ for every $s<t$. Thus, if $n/2\leq t\leq n-1$, then a set that acts transitively on the $t$-dimensional faces of an $n$-dimensional cube also acts transitively on the $(t+1)$-dimensional faces.
If $t<n-t$, then Theorem 5.7 gives $(n-t,t)\succeq(t,n-t)$ and $(n-t,t)\not\succeq(n-s,s)$ for every $s$ with $0\leq s<n-t$ except $s=t$. Thus, if $20\leq t<n/$ 2, then transitivity on the $t$-dimensional faces of an $n$-dimensional cube implies transitivity on the $(n-t)$-dimensional faces and we are in the first case again.
Similar results hold for complex regular polytopes, that is the cases $W=C_{r}\wr S_{n}$ with $r>2$ in Theorem 5.7.
6. Designs, codes and orthogonal polynomials
Certain association schemes, namely $P$- and $Q$-polynomial schemes are closely related to orthogonal polynomials in the sense that their character tables arise as evaluations of such polynomials (see [1] or [8]). The conjugacy class scheme of $C_{r}\wr S_{n}$ does not have these properties. Nevertheless, there is still a relationship to certain orthogonal polynomials, namely the Charlier polynomials.
6.1. Charlier polynomials
The Charlier polynomial of degree $k$ with parameter $a\in{\mathbb{R}}$, $a>0$, is given by
$C_{k}^{(a)}(x)=\sum_{j=0}^{k}(-1)^{k-j}\binom{k}{j}a^{-j}(x)_{j},$
where $(x)_{j}=x(x-1)\cdots(x-j+1)$ is the falling factorial. Some properties of these polynomials can be found in [15, Chapter 1.12]. The first polynomials are
$C_{0}^{(a)}(x)$ $=1,$
$C_{1}^{(a)}(x)$ $=a^{-1}x-1,$
$C_{2}^{(a)}(x)$ $=a^{-2}x^{2}-(2a^{-1}+a^{-2})x+1.$
In fact, $C^{(a)}_{1},C^{(a)}_{2},\dots$ form a system of orthogonal polynomials with respect to the Poisson distribution with mean $a$. That is, we have
(16) $\sum_{i=0}^{\infty}\frac{a^{i}}{i!}\,C^{(a)}_{k}(i)C^{(a)}_{\ell}(i)=0\quad\text{for $k\neq\ell$}.$
Sometimes the Charlier polynomials are defined to be the polynomials $a^{k}C_{k}^{(a)}$ or $(-1)^{k}C_{k}^{(a)}$. However, all of them satisfy the orthogonality relation (16).
For $g\in C_{r}\wr S_{n}$, let
$\theta(g)=\frac{1}{r}\,\left|\left\{(c,i)\in[r]\times[n]:g(c,i)=(c,i)\right\}\right|.$
That is, $\theta(g)$ is the number of fixed points of the natural action of $C_{r}\wr S_{n}$ on $[r]\times[n]$, scaled by $1/r$. Then $\theta$ is a class function of $C_{r}\wr S_{n}$ taking values in $\{0,1,\dots,n\}$ (note that $g(c,i)=(c,i)$ for some $c\in[r]$ and $i\in[n]$ implies $g(\widetilde{c},i)=(\widetilde{c},i)$ for every $\widetilde{c}\in[r]$). We write $w_{i}$ for the number of elements $g\in C_{r}\wr S_{n}$ satisfying $\theta(g)=i$. Using the inclusion-exclusion principle, one can derive the expression
$w_{i}=\frac{n!\,r^{n-i}}{i!}\sum_{j=0}^{n-i}\frac{(-1)^{j}}{j!\,r^{j}},$
see for example [6, Section 7]. We shall later see that this expression also follows from our results (see Remark 6.7).
The function $\theta$ is a class function and it defines a uniformly distributed discrete random variable on $C_{r}\wr S_{n}$. It was shown in [6, Prop. 7.4] that, for $k\leq n$, the $k$-th moment of $\theta$ equals the $k$-th moment of the Poisson distribution with mean $1/r$ (the condition $k\leq n$ was erroneously omitted in [6, Prop. 7.4]). Hence, we obtain
(17) $\mathbb{E}[\theta^{k}]=\sum_{g\in G}\frac{1}{|G|}\cdot\theta(g)^{k}=\frac{1}{|G|}\sum_{i=0}^{n}w_{i}\cdot i^{k}=e^{-\frac{1}{r}}\sum_{i=0}^{\infty}\frac{\left(\frac{1}{r}\right)^{i}}{i!}\cdot i^{k}$
for every $k\leq n$. We mention in passing that the $k$-th moment of the Poisson distribution with mean $1/r$ is given by
$\sum_{i=1}^{k}\left(\frac{1}{r}\right)^{i}S(k,i),$
where $S(k,i)$ is the Stirling number of the second kind, counting the number of partitions of a set of size $k$ into $i$ nonempty subsets.
As $C_{k}^{(1/r)}C_{l}^{(1/r)}$ is a polynomial of degree $k+l$, we find from (16) and (17) that the Charlier polynomials also satisfy the orthogonality relation
(18) $\sum_{i=0}^{n}w_{i}\,C^{(1/r)}_{k}(i)C^{(1/r)}_{\ell}(i)=0\quad\text{for $k\neq\ell$ and $k+\ell\leq n$}.$
In this section, we often use the well-known binomial transform. For a sequence $(a_{n})_{n\in{\mathbb{N}}_{0}}$, we have
(19) $s_{k}=\sum_{n=0}^{k}\binom{k}{n}a_{n}\;\Longleftrightarrow\;a_{n}=\sum_{k=0}^{n}(-1)^{n-k}\binom{n}{k}s_{k},$
also called binomial inversion, which can be obtained from the binomial theorem by straightforward calculations. Using (19), we have
(20) $(x)_{j}=a^{j}\sum_{k=0}^{j}{\binom{j}{k}}C^{(a)}_{k}(x).$
With every polynomial $f(x)=f_{n}x^{n}+\cdots+f_{1}x+f_{0}$ in ${\mathbb{R}}[x]$, we associate the class function $f(\theta)=f_{n}\theta^{n}+\cdots+f_{1}\theta+f_{0}$. This induces an algebra homomorphism from ${\mathbb{R}}[x]$ to the set of class functions of $C_{r}\wr S_{n}$. Let $\xi_{j}$ be the permutation character on ordered $j$-tuples of pairs $(c_{k},i_{k})\in[r]\times[n]$ with all $i_{k}$ pairwise distinct. By convention, $\xi_{0}$ is the trivial character of $C_{r}\wr S_{n}$. Note that
$\xi_{j}=r^{j}\,(\theta)_{j}.$
Hence, we have
(21) $C^{(1/r)}_{k}(\theta)=\sum_{j=0}^{k}(-1)^{k-j}{\binom{k}{j}}\xi_{j}\quad\text{for $k=0,1,\dots,n$},$
and from (19), we obtain
(22) $\xi_{j}=\sum_{k=0}^{j}{\binom{j}{k}}C^{(1/r)}_{k}(\theta)\quad\text{for $j=0,1,\dots,n$}.$
For $20\leq k\leq n/$ 2, we now determine the irreducible constituents of $C^{(1/r)}_{k}(\theta)$.
Theorem 6.1.
For each $k$ satisfying $20\leq k\leq n/$ 2 and each ${\underline{\lambda}}\in\Lambda_{n}(C_{r})$, we have
$\big{\langle}C^{(1/r)}_{k}(\theta),\chi^{{\underline{\lambda}}}\big{\rangle}\geq 0$
and
$\big{\langle}C^{(1/r)}_{k}(\theta),\chi^{{\underline{\lambda}}}\big{\rangle}\neq 0\;\Longleftrightarrow\;{\underline{\lambda}}(0)_{1}=n-k.$
In particular, $C^{(1/r)}_{k}(\theta)$ is a character of $C_{r}\wr S_{n}$.
Proof.
In view of (21), $C^{(1/r)}_{0}(\theta)$ is the trivial character, so the assertion of the theorem holds for $k=0$. Let $m$ be an integer satisfying $21\leq m\leq n/$ 2 and suppose that the assertion of the theorem holds for all $k$ satisfying $10\leq k\leq m-$ 1. We show that it also holds for $k=m$.
Recall that the inner product on class functions $\phi$ and $\psi$ of $C_{r}\wr S_{n}$ is given by
$\big{\langle}\phi,\psi\big{\rangle}=\frac{1}{|C_{r}\wr S_{n}|}\sum_{g\in C_{r}\wr S_{n}}\phi(g)\overline{\psi(g)}.$
It follows from the orthogonality relation (18) that
$\big{\langle}C^{(1/r)}_{k}(\theta),C^{(1/r)}_{\ell}(\theta)\big{\rangle}=0\quad\text{for $0\leq k<\ell\leq n/2$}.$
From (22), we then have
$\big{\langle}\xi_{m},C^{(1/r)}_{k}(\theta)\big{\rangle}=\binom{m}{k}\big{\langle}C^{(1/r)}_{k}(\theta),C^{(1/r)}_{k}(\theta)\big{\rangle}$
for all $k$ satisfying $20\leq k\leq n/$ 2. Since $C^{(1/r)}_{k}(\theta)$ is a character for all $k$ satisfying $10\leq k\leq m-$ 1, we find from (22) that $C^{(1/r)}_{m}(\theta)$ decomposes into those irreducible characters that occur in the decomposition of $\xi_{m}$, but not in the decomposition of $\xi_{m-1}$.
Now observe that $\xi_{\ell}$ is the induction of the trivial character of $\left(S_{1}\right)^{\ell}\times(C_{r}\wr S_{n-\ell})$ to $C_{r}\wr S_{n}$. An application of Theorem 4.5 with ${\underline{\sigma}}\in\Sigma_{n}(C_{r})$ given by ${\underline{\sigma}}(\{0\})=(1^{\ell})$ and ${\underline{\sigma}}(C_{r}\wr S_{n})=(n-\ell)$ therefore gives
$\langle\xi_{\ell},\chi^{\underline{\lambda}}\rangle\neq 0\;\Longleftrightarrow\;((1^{\ell}),n-\ell)\to{\underline{\lambda}}$
for each ${\underline{\lambda}}\in\Lambda_{n}(C_{r})$. It is easy to see that the elements ${\underline{\lambda}}$ satisfying $((1^{\ell}),n-\ell)\to{\underline{\lambda}}$ are precisely those satisfying ${\underline{\lambda}}(0)_{1}\geq n-\ell$. It follows that the irreducible constituents $\chi^{{\underline{\lambda}}}$ appearing in the decomposition of $\xi_{m}$ but not in the decomposition of $\xi_{m-1}$ are precisely those indexed by ${\underline{\lambda}}\in\Lambda_{n}(C_{r})$ with ${\underline{\lambda}}(0)=n-m$. This shows the assertion of the theorem for $k=m$. ∎
In other words, $C^{(1/r)}_{k}(\theta)$ decomposes into those characters $\chi^{{\underline{\lambda}}}$ such that the biggest part of ${\underline{\lambda}}(0)$ is $n-k$.
6.2. Designs and codes
Henceforth, we call a $((1^{t}),n-t)$-transitive set $Y$ of $C_{r}\wr S_{n}$ a $t$-design. The index of $Y$ is the number of elements in $Y$ that stabilise an arbitrary $((1^{t}),n-t)$-tabloid (so the index is the constant $c$ in the definition of a transitive set). Thus, a $t$-design is transitive on the $t$-tuples $((c_{1},i_{1}),\ldots,(c_{t},i_{t}))$, such that the elements $i_{k}\in[n]$ are pairwise distinct. We call a $t$-design of index 1 a sharply $t$-transitive set.
We also call a $((1^{n-d+1}),d-1)$-clique $Y$ of $C_{r}\wr S_{n}$ a $d$-code. Hence, for any distinct $x,y\in Y$ we have that $x^{-1}y$ does not stabilise a $((1^{n-d+1}),d-1)$-tabloid. This implies that for any distinct $x,y\in Y$ there are at least $d$ elements $i\in[n]$ such that $x$ and $y$ map $(0,i)$ to different elements.
We can also give a geometric interpretation of $t$-designs.
Example 6.2.
Let $G=C_{2}$ and $n=3$. Then $W=C_{2}\wr S_{3}$ is the symmetry group of a 3-dimensional cube. For a 1-design, we need to study the subgroup $H=C_{2}\wr S_{2}$. Geometrically, $H$ is the stabiliser of a face of the cube. Thus, $Y\subseteq C_{2}\wr S_{3}$ is a $1$-design if and only if $Y$ is transitive on the six faces of a 3-dimensional cube.
We can extend this example to $t$-designs and arbitrary $n\in{\mathbb{N}}$. Then we consider the subgroup $H=(S_{1})^{t}\times C_{2}\wr S_{n-t}=C_{2}\wr S_{n-t}$. Here, $H$ is the stabiliser of a chain of faces $F_{0}\subseteq F_{1}\subseteq\ldots\subseteq F_{t}$ of the cube where $F_{i}$ has dimension $n-t+i$. In other words, $H$ stabilises a $((1^{t}),n-t)$-flag (see Section 1). Thus, $Y\subseteq C_{2}\wr S_{n}$ is a $t$-design if and only if $Y$ is transitive on the set of $((1^{t}),n-t)$-flags of the $n$-dimensional cube. Similar interpretations can be given for other regular polytopes.
Using the results from Section 4, we can characterise codes and designs in terms of the inner and dual distribution, respectively.
Corollary 6.3.
Let $Y$ be a subset of $C_{r}\wr S_{n}$ with inner distribution $(a_{\underline{\mu}})$ and dual distribution $(a^{\prime}_{{\underline{\lambda}}})$. Then $Y$ is a $t$-design if and only if
$a^{\prime}_{{\underline{\lambda}}}=0\quad\text{for each ${\underline{\lambda}}\in\Lambda_{n}(C_{r})$ satisfying $n-t\leq{\underline{\lambda}}(0)_{1}<n$},$
and a $d$-code if and only if
$a_{{\underline{\mu}}}=0\quad\text{for each ${\underline{\mu}}\in\Lambda_{n}(C_{r})$ satisfying $n-d+1\leq{\underline{\mu}}(0)^{\prime}_{1}-{\underline{\mu}}(0)^{\prime}_{2}<n$}.$
Proof.
The first part follows from Theorem 4.6. Observe that for ${\underline{\sigma}}\in\Sigma_{n}(G)$ given by ${\underline{\sigma}}(\{0\})=(1^{t})$ and ${\underline{\sigma}}(G)=(n-t)$, we have ${\underline{\sigma}}\to{\underline{\lambda}}$ if and only if ${\underline{\lambda}}(0)_{1}\geq n-t$. This follows because in step (1) of the algorithm for the relation $\to$ in Definition 4.3, the boxes coming from ${\underline{\sigma}}(G)$ must be coloured with 00 and cannot be split up. Since the remaining $t$ boxes coming from ${\underline{\sigma}}(\{0\})$ can be coloured freely, we find that the elements ${\underline{\lambda}}\in\Lambda_{n}(G)$ satisfying ${\underline{\sigma}}\to{\underline{\lambda}}$ are precisely those satisfying ${\underline{\lambda}}(0)_{1}\geq n-t$.
The second part follows from Theorem 4.10. We have to compute the set $C({\underline{\sigma}})$ for ${\underline{\sigma}}=((1^{n-d+1}),d-1)$. First, note that the conjugacy classes of $C_{r}\wr S_{n}$ appearing in the subgroup $C_{r}\wr S_{d-1}$ are exactly the classes indexed by the set $\Lambda_{d-1}(C_{r})$. Moreover, the trivial subgroup $S_{1}$ only has the trivial conjugacy class ${\underline{\mu}}\in\Lambda_{1}(C_{r})$ with ${\underline{\mu}}(0)=1$ and ${\underline{\mu}}(g)=\emptyset$ otherwise. Hence, Theorem 4.9 gives that the elements of $C({\underline{\sigma}})$ are precisely the elements ${\underline{\mu}}\in\Lambda_{n}(C_{r})$ satisfying
${\underline{\mu}}(0)={\underline{\lambda}}(0)\cup 1^{n-d+1}\;\text{ and }\;{\underline{\mu}}(g)={\underline{\lambda}}(g)\text{ for }g\in C_{r}\setminus\{0\}$
for some element ${\underline{\lambda}}\in\Lambda_{d-1}(C_{r})$. Taking duals, we see that these are precisely the elements ${\underline{\mu}}\in\Lambda_{n}(G)$ with ${\underline{\mu}}(0)^{\prime}_{1}-{\underline{\mu}}(0)^{\prime}_{2}\geq n-d+1$. ∎
Note that $d(x,y)=n-\theta(x^{-1}y)$ defines a metric on $C_{r}\wr S_{n}$. We can therefore define codes and a distance distribution in the usual way. For a non-empty subset $Y$ of $C_{r}\wr S_{n}$, we define the distance distribution to be the tuple $(A_{i})_{0\leq i\leq n}$, where
$A_{i}=\frac{1}{|Y|}\left|\{(x,y)\in Y\times Y:d(x,y)=i\}\right|,$
and the dual distance distribution of $Y$ to be the tuple $(A^{\prime}_{k})_{0\leq k\leq n}$, where
$A^{\prime}_{k}=\sum_{i=0}^{n}C^{(1/r)}_{k}(n-i)A_{i},$
or equivalently
(23) $A^{\prime}_{k}=\frac{1}{\lvert Y\rvert}\sum_{x,y\in Y}C^{(1/r)}_{k}(\theta(x^{-1}y))=\frac{1}{\lvert Y\rvert}\sum_{x,y\in Y}C^{(1/r)}_{k}(n-d(x,y)).$
We can characterise $t$-designs in terms of zeroes in the dual distance distribution.
Proposition 6.4.
Let $Y$ be a subset of $C_{r}\wr S_{n}$ with dual distance distribution $(A^{\prime}_{k})$ and let $t$ be an integer satisfying $1\leq t\leq n$. If $Y$ is a $t$-design, then $A^{\prime}_{k}=0$ for all $k$ satisfying $1\leq k\leq t$. Moreover, the converse also holds if $t\leq n/2$. That is, if $t\leq n/2$ and $A^{\prime}_{k}=0$ for all $k$ satisfying $1\leq k\leq t$, then $Y$ is a $t$-design.
Proof.
First, assume that $Y\subseteq C_{r}\wr S_{n}$ is a $t$-design. From (23) and (21), we find that
(24) $A^{\prime}_{k}=\frac{1}{|Y|}\sum_{j=0}^{k}(-1)^{k-j}\binom{k}{j}\sum_{x,y\in Y}\xi_{j}(x^{-1}y).$
By Theorem 4.5, the permutation character $\xi_{j}$ decomposes into those irreducible characters $\chi^{{\underline{\lambda}}}$ for which ${\underline{\lambda}}(0)_{1}\geq n-j$. Moreover, since $\xi_{j}$ is a permutation character, it contains the trivial character $1_{C_{r}\wr S_{n}}$ as an irreducible constituent with multiplicity 1. From (7) and Corollary 6.3, it then follows that for all $j$ with $0\leq j\leq t$, the inner sum in (24) is
$\sum_{x,y\in Y}\xi_{j}(x^{-1}y)=\sum_{x,y\in Y}1_{C_{r}\wr S_{n}}(x^{-1}y)=|Y|^{2}.$
Hence, for all $k$ with $0\leq k\leq t$, we have
$A^{\prime}_{k}=|Y|\sum_{j=0}^{k}(-1)^{k-j}\binom{k}{j}=|Y|\delta_{k,0},$
using the binomial theorem. So we find that $A^{\prime}_{k}=0$ for all $k$ satisfying $1\leq k\leq t$.
Now, for each $k$ with $20\leq k\leq n/$ 2, we find from Theorem 6.1 that
$C_{k}^{(1/r)}(\theta)=\sum_{\begin{subarray}{c}{\underline{\lambda}}\in\Lambda_{n}(C_{r})\;:\\ {\underline{\lambda}}(0)_{1}=n-k\end{subarray}}m_{{\underline{\lambda}}}\chi^{{\underline{\lambda}}}$
for some positive integers $m_{{\underline{\lambda}}}$. Together with (23) and (7), we have that
$A^{\prime}_{k}=\frac{1}{|Y|}\sum_{\begin{subarray}{c}{\underline{\lambda}}\in\Lambda_{n}(C_{r})\;:\\ {\underline{\lambda}}(0)_{1}=n-k\end{subarray}}m_{{\underline{\lambda}}}\sum_{x,y\in Y}\chi^{{\underline{\lambda}}}(x^{-1}y)=\sum_{\begin{subarray}{c}{\underline{\lambda}}\in\Lambda_{n}(C_{r})\;:\\ {\underline{\lambda}}(0)_{1}=n-k\end{subarray}}\frac{m_{{\underline{\lambda}}}}{\deg(\chi^{{\underline{\lambda}}})}\cdot a^{\prime}_{{\underline{\lambda}}}.$
Suppose that $t\leq n/2$ and $A^{\prime}_{k}=0$ for all $k$ satisfying $1\leq k\leq t$. Since $m_{{\underline{\lambda}}}/\deg(\chi^{{\underline{\lambda}}})$ is positive, it follows that $a^{\prime}_{{\underline{\lambda}}}=0$ for all ${\underline{\lambda}}\in\Lambda_{n}(C_{r})$ with $n-t\leq{\underline{\lambda}}(0)_{1}<n$. Corollary 6.3 then implies that $Y$ is a $t$-design. ∎
We also have the following bounds.
Corollary 6.5.
Let $Y$ be a subset of $C_{r}\wr S_{n}$ and let $d$ and $t$ be the largest integers such that $Y$ is a $d$-code and a $t$-design. Then
$r^{t}\,(n)_{t}\leq\lvert Y\rvert\leq r^{n-d+1}\,(n)_{n-d+1}.$
Moreover, if equality holds in one of the bounds, then equality also holds in the other and this case happens if and only if $d=n-t+1$.
Proof.
Use Theorem 4.11 for ${\underline{\sigma}}=((1^{t}),n-t)$ and ${\underline{\sigma}}=((1^{n-d+1}),d-1)$. ∎
If we let $r=1$, then the upper bound in Corollary 6.5 becomes a well-known bound for permutation codes (see [3]). Moreover, the bounds in Corollary 6.5 can be achieved by sharply $t$-transitive sets. A construction for sharply $2$-transitive sets is given in Section 7.
The last result is about distance distributions. It turns out that the distance distribution of a subset $Y$ of $C_{r}\wr S_{n}$ is uniquely determined, provided that $Y$ is a $t$-design and a $d$-code where $d\geq n-t$.
Theorem 6.6.
Suppose that $Y$ is a $t$-design and an $(n-t)$-code in $C_{r}\wr S_{n}$. Then the distance distribution $(A_{i})$ of $Y$ satisfies
$A_{n-i}=\sum_{j=i}^{t}(-1)^{j-i}{\binom{j}{i}}{\binom{n}{j}}\bigg{(}\frac{\lvert Y\rvert}{r^{j}(n)_{j}}-1\bigg{)}$
for each $i\in\{0,1,\dots,n-1\}$.
Proof.
We have
$A^{\prime}_{k}=\sum_{i=0}^{n}C^{(1/r)}_{k}(i)A_{n-i}.$
Multiply both sides by ${\binom{j}{k}}$, sum over $k$, and use (20) to find that
$\sum_{k=0}^{j}{\binom{j}{k}}A^{\prime}_{k}=\sum_{i=0}^{n}A_{n-i}\,r^{j}\,(i)_{j}.$
Since $Y$ is an $(n-t)$-code, we have $A_{1}=\cdots=A_{n-t-1}=0$, and since $Y$ is a $t$-design, we find by Proposition 6.4 that $A^{\prime}_{1}=\cdots=A^{\prime}_{t}=0$. Moreover, we have $A_{0}=1$ and $A^{\prime}_{0}=\lvert Y\rvert$ and therefore
$\lvert Y\rvert-r^{j}\,(n)_{j}=\sum_{i=0}^{t}A_{n-i}\,r^{j}\,(i)_{j}$
for each $j\in\{1,2,\dots,t\}$. Divide by $j!\,r^{j}$ to obtain
$\sum_{i=0}^{t}{\binom{i}{j}}A_{n-i}$ $=\frac{\lvert Y\rvert-r^{j}\,(n)_{j}}{r^{j}\,j!}$
$={\binom{n}{j}}\bigg{(}\frac{\lvert Y\rvert}{r^{j}\,(n)_{j}}-1\bigg{)}$
for each $j\in\{1,2,\dots,t\}$. Now a variant of binomial inversion given by
$S_{j}=\sum_{i=0}^{t}\binom{i}{j}A_{n-i}\;\Longleftrightarrow\;A_{n-i}=\sum_{j=0}^{t}(-1)^{j-i}\binom{j}{i}S_{j}$
gives the desired result. ∎
Remark 6.7.
Note that for $Y=C_{r}\wr S_{n}$, we have that $Y$ is an $n$-design and a $1$-code. Moreover, $A_{n-i}$ equals the number of elements in $C_{r}\wr S_{n}$ with precisely $i$ fixed points, hence $A_{n-i}=w_{i}$. In this case, Theorem 6.6 gives the well-known expression
$w_{i}=\frac{n!\,r^{n-i}}{i!}\sum_{j=0}^{n-i}\frac{(-1)^{j}}{j!\,r^{j}}.$
7. Existence results
In this section, we give a simple construction of $t$-designs in $C_{r}\wr S_{n}$ using $t$-designs in $S_{n}$ and orthogonal arrays. This construction is reminiscent of constructions of designs in permutation groups using Gelfand pairs, see [2] and [14] for general groups, [19] for $S_{n}$, and [10] for $\operatorname{GL}(n,q)$. For a Gelfand pair $(G,H)$, such constructions use designs in $H$ and $G/H$ to construct designs in $G$.
It is well-known that $(C_{r}\wr S_{n},S_{n})$ is a Gelfand pair (since $C_{r}$ is abelian) whose corresponding association scheme is the Hamming scheme $H(n,r)$ (see [1, Example 2.1]). The point set of $H(n,r)$ is $(C_{r}\wr S_{n})/S_{n}$, which we identify with $C_{r}^{n}$. An orthogonal array of strength $t$, or simply a $t$-design, in $H(n,r)$ is a non-empty subset $D$ of $C_{r}^{n}$ such that the restriction of $D$ to any $t$ positions contains every element of $C_{r}^{t}$ exactly $\lvert D\rvert/r^{t}$ times. The number $\lvert D\rvert/r^{t}$ is called the index of the orthogonal array.
In what follows, we identify the elements of $C_{r}\wr S_{n}$ with pairs $(y,z)$, where $y\in C_{r}^{n}$ and $z\in S_{n}$. The following result is immediate.
Proposition 7.1.
Let $Y$ be a $t$-design in $S_{n}$ and let $D$ be an orthogonal array of strength $t$ in $H(n,r)$. Then
$\{(g,y):g\in D,y\in Y\}$
is a $t$-design in $C_{r}\wr S_{n}$.
Proof.
Note that there are constantly many elements $(g,y)$ which map a $t$-tuple $((c_{1},i_{1}),\ldots,(c_{t},i_{t}))$ with all $i_{k}$ pairwise distinct to another $t$-tuple of this type. ∎
The special case that $Y=S_{n}$ in Proposition 7.1 was also stated by Ito [14, Section 8] in a more general context.
Example 7.2.
If $r$ is a prime power and $r\geq n+1$, then it is well-known that there are orthogonal arrays of strength $t$ in $H(n,r)$ of minimum size $r^{t}$. These objects are equivalent to MDS codes in $H(n,r)$ of minimum distance $n-t+1$. Hence, if $r\geq n+1$ and there exists a $t$-design in $S_{n}$ of index $1$ (namely a sharply $t$-transitive set in $S_{n}$), then by Proposition 7.1 there exists also a $t$-design of index $1$ in $C_{r}\wr S_{n}$. For example, if $n$ is a prime power, then $\operatorname{AGL}(1,n)$ inside $S_{n}$ is a $2$-design of index $1$ and hence there exist $2$-designs of index 1 in $C_{r}\wr S_{n}$ whenever $r$ and $n$ are prime powers with $r\geq n+1$.
Now note that by Corollary 6.5, each $t$-design $Y$ in $C_{r}\wr S_{n}$ must satisfy $\lvert Y\rvert\geq r^{t}(n)_{t}$. Using the results of [16], we now show that there are always $t$-designs in $C_{r}\wr S_{n}$ almost as small as this lower bound.
Corollary 7.3.
Let $r,n,t$ be integers satisfying $r\geq 2$, $n\geq 1$, and $1\leq t\leq n$. Then there exists a $t$-design $Y$ in $C_{r}\wr S_{n}$ satisfying $\lvert Y\rvert\leq(crn)^{ct}$ for some universal constant $c>0$.
Proof.
From [16, Theorems 1.1 and 1.6], we find that there exists an orthogonal array $D$ of strength $t$ in $H(n,r)$ satisfying $\lvert D\rvert\leq\left(\frac{c_{1}rn}{t}\right)^{c_{1}t}$ and a $t$-design $Y$ in $S_{n}$ satisfying $\lvert Y\rvert\leq(c_{2}n)^{c_{2}t}$ for some universal constants $c_{1},c_{2}>0$. The statement of the corollary then follows from Proposition 7.1. ∎
Note that Corollary 7.3 also gives the existence of more general designs in $C_{r}\wr S_{n}$, since in view of Theorem 5.7, every $t$-design is also a $(\sigma,\rho)$-design, where $(\sigma,\rho)$ is a double partition of $n$ such that $\rho_{1}\geq n-t$.
8. Application to projective planes
In this section, we explore a connection to projective planes. For an introduction, we refer the reader to [12]. We make use of two well-known results on the existence of projective planes.
Theorem 8.1 (Bose [4]).
There exists a projective plane of order $n$ if and only if there exists a complete set of $n-1$ mutually orthogonal latin squares of order $n$.
It is well-known that there exists a complete set of mutually orthogonal latin squares of order $n$ if and only if there exists an orthogonal array of strength 2 with $n+1$ columns, $n$ symbols and index 1. Thus, the existence of a projective plane of order $n$ is equivalent to the existence of a $2$-design in the Hamming scheme $H(n+1,n)$ of index 1.
It is also well-known that the existence of a finite projective plane of order $n$ is connected to the existence of a sharply 2-transitive set in $S_{n}$, so a 2-design in $S_{n}$ of index 1.
Theorem 8.2 (Hall [11]).
There exists a projective plane of order $n$ if and only if there exists a sharply 2-transitive subset of $S_{n}$.
Using Proposition 7.1, we can construct a $2$-design in $C_{r}\wr S_{n}$ from a $2$-design in $S_{n}$ and a $2$-design in $H(n,r)$. This gives the following theorem.
Theorem 8.3.
Let $n$ be a positive integer. If there is no $2$-design of index 1 in $C_{n-1}\wr S_{n}$, then there is no projective plane of order $n-1$ or there is no projective plane of order $n$.
Proof.
We show that the existence of a projective plane of order $n-1$ and $n$ implies the existence of a $2$-design of index 1 in $C_{n-1}\wr S_{n}$.
A projective plane of order $n$ gives rise to a 2-design in $S_{n}$ of index 1. A projective plane of order $n-1$ gives rise to a 2-design in $H(n,n-1)$ of index 1. Now Proposition 7.1 gives a $2$-design in $C_{n-1}\wr S_{n}$ of index 1. ∎
If $n$ is a prime power, then there exists a projective plane of order or $n$. In this case, the non-existence of a 2-design in $C_{n-1}\wr S_{n}$ implies the non-existence of a projective plane of order $n-1$. Similarly, if $n-1$ is a prime power, then the non-existence of a 2-design in $C_{n-1}\wr S_{n}$ implies the non-existence of a projective plane of order $n$.
Note that it is possible that there exist 2-designs of index 1 in $C_{n-1}\wr S_{n}$ that do not come from the construction of Proposition 7.1. In this case, Theorem 8.3 does not apply. More detailed analysis of the structure of transitive sets in $C_{n-1}\wr S_{n}$ is needed to explore this connection further.
Acknowledgements
The first author would like to thank Barbara Baumeister, Alena Ernst, Karen Meagher and Charlene Weiß for helpful discussions and proofreading of this work. Sadly, the second author passed away during the writing process. The first author is grateful to him for the supervision and the fruitful collaboration on this article.